### v1

# TASER: Table Agents for Schema‑guided Extraction and Recommendation

Abstract
Real-world financial documents report essential information about an entity’s financial holdings that can span millions of different financial instrument types. Yet, these details are often buried in messy, multi-page, fragmented tables - for example, 99.4% of the tables in our dataset have no bounding boxes with the maximum number of rows amounting to 426 per table across 44 pages. To tackle these unique challenges from real-world tables, we present a continuously learning, agentic table extraction system, TASER (Table Agents for Schema-guided Extraction and Recommendation) that extracts highly unstructured, multi-page, heterogeneous tables into normalized, schema-conforming outputs. Our table agents execute on table detection, classification, extraction, and recommendations by leveraging an initial schema. Then, our Recommender Agent reviews the outputs, recommends schema revisions, and decides on the final recommendations, enabling TASER to outperform existing table detection models such as Table Transformer by 10.1%. Within this continuous learning process, we highlight that larger batch sizes result in a 104.3% increase in schema recommendations that are actionable and utilized, resulting in a 9.8% increase in extracted holdings - highlighting the importance of a continuous learning process. To train TASER, we have manually labeled 22,584 pages (28,150,449 tokens), 3,213 tables for \$731,685,511,687 of holdings culminating in one of the first real financial table datasets. We release our dataset TASERTab to enable the research community to access real-world financial tables and outputs. Our results highlight the promise of agentic, schema-guided extraction systems for robust understanding of real-world financial tables.
Figure 1: Complexity of Holdings Table in Regulatory Filings. In the original format, multiple data attributes are displayed in a single line, with no bounding boxes, rendering the generation of structured outputs highly challenging. TASER enables the generation of structured outputs from highly variable, multi-page financial tables for complex instrument holdings. Negative quantities or market values denote short positions. See Appendix L for additional outputs.
Figure 2: Variety and complexity of financial tables. From leftmost column - for a single financial filing type, such as annual reports, there are a variety of formats by which these reports are filed. Within each filing, there are numerous table types - spanning from Performance, Financial Holdings, or Cash Flow Tables. Each table taxonomy houses very different types of information and has different objectives. For example, the Cash Flow Table houses information relating to operational cash movements for an entity. In contrast, the Financial Holdings Table (which is our table of interest in this paper) displays all financial instruments that an entity holds. For the Financial Holdings Table, there are numerous layout structures and formats by which tables appear, as seen in the rightmost column. Due to the extreme heterogeneity of formatting, document layout, and table structure, traditional table extraction methods fail to perform for financial filings.
Model Modality Primary Task Promptable
Camelot Vision + Spatial Heuristic Table Detection & Parsing No
Table Transformer Vision Detection & Structure Recognition No
TaPas Text Table‑based QA Partially
TAPEX Text Programmatic Extraction (SQL) Partially
TASER (ours) Vision + Text Schema‑guided Extraction Yes
Table 1: Comparison of representative table extraction and reasoning models. Our work extends prior methods by introducing a fully agentic, schema-guided extraction framework for highly complex financial tables, leveraging prompt-based self-refinement and continuous schema adaptation.
1 Introduction
Financial documents, particularly annual regulatory filings for funds, house tables that govern \$68.9 trillion of investments globally (Investment Company Institute 2024). In perspective, \$68.9 trillion is more than twice the total Gross Domestic Product (GDP) of the United States (\$29.1 trillion) (WorldBank 2025). This critical data is housed in the Financial Holdings Table (Figure 1), which outlines the entirety of an entity’s investment holdings (U.S. Congress 1934; EU Commission 2019); this table is the longest table in terms of row count (with a maximum row count of 426 rows) - more than double the average row count of all other table types (Table 7). These Financial Holdings Tables are not only very long but also highly heterogeneous in layout structure and visually complex (Figure 2). While generating structured outputs from these tables is critical for many regulatory and financial institutions (Cho et al. 2024), there is a relative dearth of studies that focus on understanding the complexity of Financial Holdings Tables, compared to web or SQL tables (Herzig et al. 2020; Pasupat and Liang 2015; Zhong, Xiong, and Socher 2017). Therefore, the following challenges exist in terms of parsing Financial Holding Tables into structured, machine-readable outputs (Figure 2):
[…]
    One-to-many relationships between a document and the tables it houses exacerbates standard model performance for table detection or structure recognition tasks.
[…]
    Financial Holdings Tables span across multiple pages, rendering models that operate at the page-level inefficient.
[…]
    Financial instruments are highly complex, with embedded structures and nested hierarchies. Therefore, details are often clumped in a single cell as seen in Figure 1.
[…]
    Tabular layouts are heterogeneous with no bounding boxes, mixing tables, text blocks, footnotes, and images, often without consistent labeling or alignment. Moreover, we identify that 99.4% of tables in our dataset, TASERTab, lack bounding boxes or gridlines to efficiently identify a single cell.
These challenges drive our motivation to construct our agentic table extraction methodology that is capable of goal-driven parsing and self-refinement, with the ability to decide on the optimal schema, continuously learning and reasoning from errors.
Contribution 1: We propose a novel table extraction methodology, TASER (Table Agents for Structured Extractions and Recommendation) that undertakes table detection, classification, extraction, and recommendations in a singular pipeline by leveraging the provided schema. We compare our methodology against predominant table extraction methodologies and report TASER’s 10.1% outperformance against Table Transformer (Yang et al. 2022) for detection.
Contribution 2: Once the initial schema is given, we prove the effectiveness of our Recommender Agent to continuously improve the initial schema - reflecting a tunable and continuous self-learning loop. Throughout our training, we discover that small batches are optimal in providing diverse and comprehensive recommendations to the original schema–however, at the cost of redundant recommendations. In contrast, large batches drive high precision recommendations - however, at the cost of diversity. Thus, our results establish that self-learning via agents for table extraction is tunable; through adjusting batch size, we can directly control schema refinement to maximize actionable coverage while minimizing redundancy.
Contribution 3: We have constructed a manually labeled dataset TASERTab of ground truth labels for 3213 real-world Financial Holdings Tables amounting to \$731.7B in value. We sourced the filings directly from fund websites, labeled the total net assets for each fund, and recorded the span of each Financial Holdings Table. We believe that this is the first dataset of its kind for the research community to access real-world financial tables with structured outputs.
2 Related Work
Information & Table Extraction: Early information extraction relied on statistical models (HMMs (Borkar, Deshmukh, and Sarawagi 2001), CRFs (Lafferty, McCallum, and Pereira 2001), heuristics (Press 2003), and graph-based layouts (Liu et al. 2019; Qian et al. 2019; Meuschke et al. 2023), but still struggle with complex, heterogeneous tables.
Table Representation Learning: Transformer-based table understanding and QA include TaPaS (Herzig et al. 2020), TaBERT (Yin et al. 2020), TaPEX (Liu et al. 2022), TURL (Deng et al. 2020), TUTA (Wang et al. 2021), and TableFormer (Yang et al. 2022). These methods encode text, structure, and layout, but few are benchmarked on long, dense, multi-page financial reports.
LLMs for Structured Data: General LLMs have strong performance for schema-conformant extraction via fine-tuning & prompting (Brown et al. 2020; Ouyang et al. 2022; Paolini et al. 2021; Wei et al. 2023; Lu et al. 2022; Liu and Contributors 2024), while multimodal approaches (LayoutLM (Xu et al. 2020, 2021; Huang et al. 2022), DONUT (Kim et al. 2022), DocFormer (Appalaraju et al. 2021), UniTable (Peng et al. 2024), and Table Transformer (Smock, Pesala, and Abraham 2021; Carion et al. 2020) improve layout awareness but still lag on long, fragmented tables (Zhao et al. 2024).
Financial Document Parsing: Includes multimodal extraction (Watson and Liu 2020), expert agent pipelines (Cho et al. 2024), and large-scale benchmarks (DocILE (Šimsa et al. 2023), BuDDIE (Wang et al. 2025), Kleister (Stanisławek et al. 2021), FinTabNet (Zheng et al. 2020)).
Agentic and Recursive Extraction: Recent methods cast LLMs as agents capable of iterative extraction and self-correction (Shen et al. 2023; Roucher et al. 2025; Yao et al. 2023; Shinn et al. 2023; Watson et al. 2023; Yuan and Xie 2025). Prompt-based feedback, introspective refinement, and episodic memory frameworks (Madaan et al. 2023; Shinn et al. 2023; Yao et al. 2023; Wei et al. 2023; Wang et al. 2023; Shridhar et al. 2023; Du et al. 2024; Lee et al. 2024; Lewis et al. 2021; Schulhoff et al. 2025) drive improvements in reasoning for complex extraction.
LLM-Augmented Aggregation: LLMs are increasingly used for hierarchical and agglomerative clustering, enhancing interpretability and item-set quality (Viswanathan et al. 2024; Pattnaik et al. 2024; Zhang, Wang, and Shang 2023) by leveraging semantic reasoning in merge decisions.
3 Methodology
3.1 System Architecture
Our methodology is composed of three core Large Language Model (LLM) agents, each with a distinct role. We conduct rigorous ablations to evaluate the importance of each agent.
  1. 1.
    Detector Agent: Identifies candidate pages containing Financial Holdings Tables leveraging the initial schema provided. The prompt is tuned to maximize recall as we do not want to miss any Financial Holdings Tables. We provide our prompts in the Appendix (Figure 12).
  2. 2.
    Extractor Agent: Processes detected pages by prompting the LLM with the current Portfolio schema embedded in the prompt context. The LLM’s output is validated inline against the schema using Pydantic & Instructor, producing a set of structured, type-checked instrument entries (Figure 1).
  3. 3.
    Recommender Agent: Reviews unmatched extractions containing both false and true positives. A false positive is a spurious extraction (e.g., headers/subtotals/footnotes/OCR noise or cells from non-holdings tables) that fails schema/type/consistency checks; a true positive is a valid holdings field from a genuine row that the current schema cannot yet classify (e.g., new security type or column alias) but passes those checks. The agent first filters false positives by re-validating each candidate under the current schema; it proposes schema modifications for the remaining true positives and, per class, recommends the minimal change that correctly classifies each item.
All agents interact through explicit artifacts: structured outputs, episodic error stacks, and schema definitions. Output validation is integrated into each agent’s forward pass via instructor (Liu and Contributors 2024). TASER implements a recursive feedback loop, where errors and unmatched holdings identified in the initial extraction are escalated to the Recommender Agent, which provides recommendations to refine the schema and triggers re-extraction. This loop repeats until all entries are either matched or a stopping criterion is met (no new schema modifications proposed). The pipeline is fully parallelizable for multi-page and multi-entity filings. A schematic of the full agentic pipeline is shown in Figure 3.
Figure 3: Agentic feedback loop for schema-guided table extraction and continuous schema recommendations. The pipeline begins with raw document ingestion, where candidate pages are evaluated by a Table Detector Agent. Only pages identified as Financial Holdings Tables proceed to the extraction stage, where our Extractor Agent parses each table into a structured, schema-conformant response (e.g., extracting option instrument fields as shown). In the Refinement stage, outputs are partitioned into declared (matched) and unmatched holdings. Unmatched holdings are grouped into batches, which trigger a schema recommendation and feedback loop from the Recommender Agent (Figure 4). The Recommender Agent proposes schema refinements, reconciles its recommendations, and finally, updates the schema, enabling TASER to adapt and improve extraction fidelity over successive iterations.
Figure 4: Schema-Guided Agentic Extraction and Refinement Loop. The extraction pipeline begins with an Initial Schema Definition (v1), which guides the LLM Extractor Agent as it processes the raw Holdings Table to produce Declared Holdings. Holdings that do not match the schema are routed as Unmatched Holdings, triggering the generation of Schema Update Suggestions. These suggestions are reviewed, clustered, and aggregated by our Validation Agent before updating the schema (v2), replacing the prior definition and closing the agentic feedback loop. This process enables continuous schema refinement and robust extraction from heterogeneous, visually complex financial tables.
3.2 Initial Schema Definition and Application
TASER’s extraction process is anchored by an explicit, user-modifiable Portfolio schema that defines the target structure for Financial Holdings Tables. We implement this schema using Pydantic models; our initial schema reflects our current knowledge by leveraging external knowledge (U.S. Congress 1934). Each schema consists of a base Instrument model, subclassed for common asset types (e.g., Equity, Bond, Option, Swap, Forward, Future, Debt, EquityLinkedNote) and an Other class for uncategorized assets. Each subclass specifies instrument-specific fields and validation logic (see App. H).
Schema-Guided Extraction: For each candidate page, the Extractor Agent prompts the LLM with the current schema embedded in the prompt context. The LLM is instructed to return a structured output, which is immediately parsed and validated against the schema using Pydantic’s type checking and validation logic. Outputs that fail schema validation (e.g., missing fields, type errors, or undeclared instruments) are flagged as feedback for re-extraction.
Schema Recommendations for Iterative Refinement: We formalize schema refinement as an iterative, LLM-driven clustering process that updates the original schema to accommodate unmatched holdings and novel instrument types discovered during extraction (Novikov et al. 2025; Pattnaik et al. 2024; Zhang, Wang, and Shang 2023). At each iteration, the agent operates on the episodic error stack to propose schema modifications, and extraction is retried using the updated schema. This process continues until all entries are matched or no further improvements are possible. Let $H=\{h_{1},h_{2},\ldots,h_{N}\}$ denote the set of unmatched holdings, and let $\Sigma^{(0)}$ be the initial schema. For each iteration $\ell$, we define:
[…]
    $H^{(\ell)}$: Unmatched holdings at iteration $\ell$.
[…]
    $\Sigma^{(\ell)}$: Current schema.
[…]
    $g_{\theta}$: LLM-based schema suggestion function.
[…]
    $B$: Batch size for error grouping.
The refinement loop (Algo. 1) proceeds as follows:
  1. 1.
    Partition $H^{(\ell)}$ into batches of size at most $B$.
  2. 2.
    For each batch, invoke $g_{\theta}$ with batch errors and $\Sigma^{(\ell)}$ to propose schema modifications.
  3. 3.
    Aggregate all proposals, cluster/reconcile if necessary, and select the modifications to apply.
  4. 4.
    Update schema to $\Sigma^{(\ell+1)}$ and re-extract.
  5. 5.
    Update error stack and repeat until $H^{(\ell+1)}$ is empty or no new schema changes are suggested.
Algorithm 1 LLM Iterative Schema Refinement
0: Unmatched holdings $H=\{h_{1},h_{2},\ldots,h_{N}\}$, initial schema $\Sigma^{(0)}$, LLM schema suggestion function $g_{\theta}$, batch size $B$, stopping criterion $T$
1: Initialize $\ell\leftarrow 0$
[…]
4:while not stopping criterion $T$ met do
5:  Partition $H^{(\ell)}$ into batches $H^{(\ell)}_{j}$ of size at most $B$
6:  $S^{(\ell)}\leftarrow\emptyset$ {Suggested schema modifications}
7:  for each batch $H^{(\ell)}_{j}$ do
8:   $S_{j}^{(\ell)}\leftarrow g_{\theta}(H^{(\ell)}_{j},\Sigma^{(\ell)})$
9:   $S^{(\ell)}\leftarrow S^{(\ell)}\cup S_{j}^{(\ell)}$
10:  end for
11:  $S_{\text{selected}}^{(\ell)}\leftarrow\text{AggregateAndSelect}(S^{(\ell)})$ {Aggregate suggestions}
12:  $\Sigma^{(\ell+1)}\leftarrow\text{UpdateSchema}(\Sigma^{(\ell)},S_{\text{selected}}^{(\ell)})$
[…]
14:  if $H^{(\ell+1)}=\emptyset$ then
15:   break
16:  end if
17:  $\ell\leftarrow\ell+1$
18:end while
19:return $\Sigma^{(\ell+1)}$
3.3 Ablation Strategies and Efficiency
We conduct a systematic ablation study to isolate the contributions of schema-guided extraction, prompt engineering, and agentic feedback within TASER. Four main extraction strategies are compared:
  1. 1.
    Raw Text Prompting: The LLM is prompted only with the page text; extraction is based solely on a yes/no detection.
  2. 2.
    Structured Chain-of-Thought (CoT): Prompts include a minimal schema and few-shot examples, eliciting explicit reasoning traces before a final boolean decision.
  3. 3.
    Full Schema Prompting: The full Portfolio schema is embedded in the prompt, instructing the LLM to return structured, schema-conformant entries.
  4. 4.
    Direct Schema Application: The schema is directly applied to parsed page content without prior detection; extraction succeeds if any schema sub-model instantiates.
Table 2 reports detection and extraction accuracy for each strategy. Full Schema Prompting achieves the highest precision, F1, and accuracy, while maintaining perfect recall. Table 3 summarizes extraction fidelity in terms of total absolute dollar difference and percent unaccounted value, and Table 10 compares computational efficiency in terms of tokens and latency.
Method Recall (%) Precision (%) F1 (%) Accuracy (%)
Camelot (Hybrid) 56.92 23.46 33.23 47.35
Table Transformer (Smock, Pesala, and Abraham 2021) 99.76 32.75 49.31 46.27
TASER
(a) Raw Text Prompting 100.0 38.62 55.73 58.38
(b) Structured CoT 100.0 34.42 51.21 50.10
(c) Full Schema Prompting 100.0 43.43 59.44 66.35
(d) Direct Schema Application 100.0 41.84 58.30 63.99
Table 2: Detection performance across ablation strategies. While all TASER ablations achieve perfect recall (100%), Full Schema Prompting yields the highest precision (43.43%), F1 score (59.44%), and overall accuracy (66.35%), underscoring the value of embedding the complete Portfolio schema in the detection prompt. See Table 4 for all Camelot configurations.
Method Total Absolute Difference (USD) Unaccounted (% of total)
TASER
(a) Raw Text Prompting \$ 107,066,845 0.015%
(b) Structured CoT \$ 120,577,458 0.016%
(c) Full Schema Prompting \$ 102,836,797 0.014%
(d) Direct Schema Application \$ 118,881,312 0.016%
Table 3: Extraction metrics for each ablation strategy. We report the total absolute difference (TAD) between extracted and ground‐truth net asset values and the percentage of total holdings unaccounted for out of \$731.7 billion. Lower TAD and unaccounted percentages indicate higher dollar‐value fidelity; Full Schema Prompting achieves the best performance (TAD = \$102.8M, 0.014% unaccounted), demonstrating superior accuracy in financial extraction.
4 Experimental Setup
All experiments use gpt-4o-2024-11-20 as the LLM.
4.1 Evaluation Metrics
Detection Metrics: We report recall, precision, F1, and accuracy for table detection, prioritizing recall to eliminate the risk of missed financial holdings tables in extraction.
Extraction Metrics: We also assess extraction completeness by comparing TASER’s outputs to the ground truth labels. Each Financial Holdings Table houses a total net asset value, which we manual label for . We compare our labeled ground truth with the addition of all our extractions - this will be dubbed the total absolute difference (TAD).
Schema Refinement Metrics: Coverage is the fraction of unmatched holdings aligned with at least one schema suggestion, using RapidFuzz string similarity with a lenient ($\geq 70$) threshold. We also report the number of new matched holdings after re-extraction with the suggested schemas added to Portfolio. Diversity is the average pairwise Levenshtein distance between suggestion attributes (name and generated schema). Collision rate denotes the proportion of duplicate schemas among suggestions (reported on name and generated schema).
Dataset: We curate a diverse corpus of financial filings totaling 22,584 pages, 28M tokens, and \$731.7B in holdings. Among 3,213 tables, 57.53% exhibit hierarchical structure (via spanning cells). All Portfolio of Investment tables (100%) are hierarchical. While 39% of portfolios are single-page, 60.2% span multiple pages. The average length is 3.24 pages ($\sigma=3.41$, max = 19). This variability underscores the need for multi-page detection and consolidation.
Figure 5: Left: Cumulative unique schemas per iteration; larger batches discover schemas rapidly but plateau quickly. Right: Cumulative unique schemas per unmatched holding seen; smaller batches ultimately yield more unique schemas but require more suggestions and generate more redundancy.
5 Results and Discussion
5.1 Quantitative Evaluation
Detection: Table 2 shows that all TASER ablations achieve perfect recall ($\sim$100%), but precision ranges from 32.8% (Table Transformer) up to 43.4% (Full Schema Prompting), driving F1 scores between 49.3% and 59.4%. Embedding the full Portfolio schema in the prompt boosts precision by over 10% relative to the vision‑only baseline and yields the highest F1 (59.4%) and accuracy (66.4%), demonstrating that in‑context schema guidance is critical for filtering true holdings tables from noisy pages.
Extraction: Table 3 confirms that schema anchored extraction improves dollar‑value fidelity. Full Schema Prompting attains the lowest total absolute difference (\$102.8M) and the smallest unaccounted share (0.014%), outperforming Raw Text Prompting (\$107.1M, 0.015%) and Structured CoT (\$120.6M, 0.016%). Direct Schema Application (skipping detection) incurs a higher error (\$118.9M; 0.016%) by parsing spurious non‑holding pages. Therefore, full-schema detection followed by schema-driven extraction filters to true holdings tables, tightening dollar-value estimates.
5.2 Success Highlights
Cross-Document Consistency: TASER classifies and extracts holdings tables with varying titles (e.g., ”Portfolio of Investments”, ”Schedule of Holdings”, or ”Investment Portfolio”) and diverse structural formats. Despite the immense complexity of inputs, TASER consistently extracts and transforms the tables, ensuring that the final output appears as if sourced from a uniform set.
Contextual Understanding: TASER excels in handling contextual nuances within financial documents, such as interpreting negative values denoted by parentheses (e.g., (140)) in zero-shot settings. Such domain-specific attributes are important in understanding financial tables.
Extracting Intricate Semantics: TASER demonstrates a strong semantic understanding of financial terminology, which empowers it to not only extract but also correctly comprehend the financial information it processes. For instance, TASER adeptly parsed the table entry “GBP 4,700,000 — UK Treasury 0% 19/02/2024 — 4668 — 1.48,” correctly identifying the holding as a bond and extracting its attributes: quantity, market value, coupon rate, maturity date, and issuer.
Figure 6: Reduction in total absolute difference (TAD) after resolving unmatched holdings. Remaining TAD is calculated after sequential reconciliation of unmatched holdings. Most error reduction is achieved by resolving the most significant unmatched holdings; additional reconciliation yields only incremental improvement, consistent with a heavy-tailed value distribution (Figs. 8 & 9).
5.3 Batch Size Tradeoffs in Schema Refinement
Figure 5 (left) reveals that larger batch sizes (250, 500) rapidly expand the schema set in the first few iterations, achieving faster initial coverage as measured by iterations. However, this early acceleration comes at the cost of quickly reaching saturation, after which few new unique schemas are discovered. In contrast, smaller batch sizes (10, 50, 100) require more iterations to reach the same number of unmatched holdings seen, but continue yielding new unique schemas for much longer, resulting in the highest overall diversity when normalized by data processed (Figure 5, right). This improvement in coverage for smaller batches is offset by increased redundancy. As shown in Appendix 7, smaller batches incur substantially more duplicate or overlapping schema suggestions, reflecting a more granular and exploratory nature of their schema search. Overall, these results highlight a key tradeoff in agentic schema refinement: larger batches accelerate early-stage discovery but plateau quickly, while smaller batches maximize cumulative schema diversity at the cost of higher redundancy and computation. Our results establish that schema refinement via agentic suggestion and feedback is both tractable and tunable: by adjusting batch size, practitioners can directly control the balance between coverage and efficiency, guiding schema evolution to maximize actionable suggestions while minimizing redundancy and annotation waste. This provides a principled, quantitative framework for automated schema refinement in noisy, open-ended domains
Schema Diversity and Utilization: Schema diversity, as measured by the average pairwise Levenshtein distance, is maximized for moderate batch sizes (100–250), as shown in Appendix K. While larger batch sizes (500) yield a higher proportion of utilized schemas—up to 59%—smaller, more diverse batches tend to have lower utilization rates (Table 9). Furthermore, the accretive gain in 402 additional unique schemas yielded only marginal improvements in holding coverage (6.1%). Figure 10 illustrates this tradeoff: smaller batch sizes cover more unmatched holdings at the expense of efficiency (96.1% coverage for 29.0% utilization at batch size 10), whereas larger batches achieve higher schema utilization (59.0% at batch size 500), a greater proportion of actionable suggestions. This spectrum allows TASER to target either exhaustive coverage or more efficient, high-value schema evolution, depending on application needs.
5.4 Schema Refinement Optimality
Systematic ablation of batch size during schema refinement for unmatched financial holdings demonstrates a tunable tradeoff between coverage—the absolute number of holdings addressed—and utilization—the efficiency of actionable schema suggestion. By selecting appropriate batch sizes, schema refinement can be tailored to maximize coverage, efficiency, or a balanced combination, supporting principled, automated schema evolution in real-world information extraction. Practically, these competing objectives can be summarized as follows: small batches maximize recall, ensuring that most unmatched entities are eventually addressed, while large batches drive high-precision, low-waste schema discovery. Practically, this means that a mixed or adaptive batching strategy may be optimal—using larger batches to quickly identify high-yield schemas, followed by smaller, targeted batches to exhaustively cover residual unmatched cases.
Improvements in TAD: Resolving the largest unmatched holdings yields a reduction in total absolute difference (TAD) of approximately 7–10% across batch sizes, with the majority of improvement achieved by reconciling just the top 10–20% of holdings (Table 5). Beyond this, further reconciliation produces only marginal gains, reflecting the highly heavy-tailed nature of recoverable value.
6 Conclusion
By combining LLM agents with dynamic, schema‑anchored prompting and recursive validation loops, TASER offers a new proposal for extracting and normalizing complex holdings tables from raw financial filings. TASER’s high precision and recall across diverse financial layouts underscore the potential of agentic, domain‑aware reasoning for scalable and accurate document understanding.
Appendix A Disclaimer
This paper was prepared for informational purposes by the Artificial Intelligence Research group of JPMorgan Chase & Co. and its affiliates (”JPMorgan”) and is not a product of the Research Department of JPMorgan. JPMorgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
Appendix B Limitations
Despite its strong performance, TASER remains susceptible to errors in low-resolution or scanned PDFs, where visual degradation can hinder accurate extraction. Ambiguities in financial documents, such as undefined asset classes or implicit references, pose challenges that cannot always be resolved without external knowledge or manual intervention. While recursive prompting enhances completeness, it introduces added latency and computational overhead. Additionally, TASER relies on prompt-based weak supervision due to the lack of fine-grained, labeled datasets for complex instrument types, which may limit generalization. Finally, TASER does not yet model interactions between table rows or instrument relationships beyond the schema level, which may affect downstream tasks such as portfolio risk analysis or exposure aggregation.
Appendix C Camelot Table Parsing Modes
For completeness, we also compare TASER’s detection performance against the four table detection modes in Camelot [^1] https://github.com/camelot-dev/camelot. The best-performing variant (Hybrid) achieves an F1 score of 0.33, still below TASER’s weakest ablation (0.51). Full results are presented in Table 4. Note that our financial tables primarily consist of unruled, whitespace-separated tables with alignment-based structure. Below is a brief summary of each mode:
[…]
    Stream: Groups text using whitespace and y-axis alignment. Suitable for unruled tables, but yielded low precision on our data ($F1=21.6\%$).
[…]
    Lattice: Uses image-based line detection to extract ruled tables. Less effective for our dataset due to the rarity of bordered layouts ($F1=13.3\%$).
[…]
    Network: Detects tables via text alignment patterns using bounding boxes. Performs better on our format, which lacks explicit ruling ($F1=18.6\%$).
[…]
    Hybrid: Combines Network’s structure with Lattice’s grid refinement. Achieved the highest F1 score (33.23%) among Camelot modes, confirming the benefit of integrating both visual and alignment cues.
Figure 7: Cumulative collisions per unmatched holding; smaller batches incur more collisions, reflecting greater redundancy.
Appendix D TASER Annotation Process
We manually sourced each financial document directly from the fund entity’s public website, ensuring broad coverage across instrument types. Annotations were performed at the page, table, and holdings level (which may span hundreds of pages). For every filing and fund, we recorded the page-span for the portfolio of investments table and the net asset value across all holdings for that fund.
Appendix E TASER Dataset Release
TASER is built on public fund documents. Our release will include labels for the positions of holdings tables, the recorded net asset value, the fund name, multi-page spans, and a URL reference to the public fund document. Each pdf filing is hosted by the fund’s advisor, as required by regulation.
Batch Remaining TAD Reduction NAV
Size TAD (\$) (%) Extracted (\$)
500 94,843,638 7.8% 7,993,158
250 94,185,693 8.4% 8,651,103
100 95,985,588 6.7% 7,851,209
50 92,781,421 9.8% 10,025,376
10 93,032,549 9.6% 9,804,248
Table 5: Remaining Total Absolute Difference (TAD, \$) and Net Asset Value (NAV, \$) extracted from reconciled unmatched holdings by batch size.
Appendix F Document Preprocessing
For each PDF filing, TASER extracts raw text, layout metadata, and embedded images using a hybrid pipeline based on pdfplumber. Each page is parsed into normalized text blocks and layout primitives, preserving spatial relationships and read order. Minimal normalization is applied, including Unicode cleanup and header/footer removal. Each page object includes:
  • Raw text blocks (reading order preserved)
  • Bounding boxes and font metadata
  • Embedded images (if any)
We apply Unicode normalization (NFKC), whitespace collapse, and filter out repeated headers/footers via regex matching. Optionally, OCR is performed if text extraction fails. Code and parameters are available upon request.
Figure 8: Heavy-tailed distribution of value recovery from unmatched holdings across batch sizes. We report the Lorenz curves for the cumulative fraction of value recovered as a function of the fraction of “other” holdings resolved. For all batch sizes, a small number of matches account for the vast majority of recovered net asset value, while most resolved holdings contribute negligibly. The bow of each curve away from the diagonal illustrates the extreme concentration of recoverable value in the “head,” characteristic of a heavy-tailed regime.
Figure 9: Cumulative recovery fraction vs. number of holdings resolved. Cumulative fraction of total value recovered as a function of the number of unmatched holdings resolved (log-log scale). The steep initial rise for each batch size indicates that the largest recoveries are concentrated among the first few resolved holdings; subsequently, improvement plateaus, indicating diminishing returns from resolving additional holdings in the long tail.
Appendix G Parallelization and Fund Construction
Extraction: To efficiently process large, multi-page filings, TASER employs parallelization (20 workers) at both the document and page levels. Each agent operates asynchronously across document batches: Detector and Extractor agents process candidate pages in parallel, while the Recommender agent operates downstream on the resulting artifacts.
Merging: For fund-level construction, extracted tables from consecutive pages are merged deterministically. Entity resolution is performed by matching predicted fund names and table headings across pages, while units and currencies are normalized to a consistent reporting standard through a boolean flag value_in_thousands. Partial extractions are reconciled using strict types in the response model, whose validation errors re-prompt the LLM on specific extraction errors to ensure a unified, schema-conformant portfolio representation for each fund.
Dataset Avg. # Tables Avg. Rows Avg. Columns Avg. Spanning
Per Topology Per Table Per Table Cells per Table
SciTSR 5.70 9.28 5.19 0.77
PubTabNet 4.13 14.05 5.39 2.24
FinTabNet 11.80 11.93 4.36 1.01
PubTables-1M 3.78 13.41 5.46 3.01
TASERTab 11.00 53.70 6.36 2.67
Table 6: Complexity of table instances across datasets. TASERTab exhibits almost five times the number of rows compared to other datasets. The maximum row count in TASERTab is 426 rows across 44 pages for a single Financial Holdings Table.
Dataset # Unique Cell Avg. # Tables Avg. Rows Avg. Columns Maximum Page Span
# Tables Topologies Per Topology Per Table Per Table
Financial Holdings Table 1933 621 3.11 53.7 6.36 44
All Other Tables 1280 331 4.32 26.9 3.87 37
Table 7: Complexity of Financial Holdings Tables
Appendix H Schema Definitions and Portfolio Model
Portfolio Base Model:
The core Instrument base model in our Pydantic model is subclassed into the following classes (see Figure 13 for the full class diagram). This is our initial schema composed of some of the most well-known financial instruments:
[…]
    Equity: a share of ownership in a corporation, representing residual claims on earnings and assets.
[…]
    Bond: a fixed‐income security issued by governments or corporations, paying periodic coupons and returning principal at maturity.
[…]
    Future: an exchange‐traded contract obligating the buyer or seller to transact an asset at a predetermined price on a specified future date.
[…]
    Forward: an over‐the‐counter agreement to buy or sell an underlying asset at a set price on a future date, customizable but counterparty‐risky.
[…]
    Swap: a bilateral contract to exchange cash flows (e.g., fixed vs. floating interest rates or different currencies), with terms set at initiation.
[…]
    Option: a derivative granting the right, but not the obligation, to buy (call) or sell (put) an underlying asset at a specified strike price before or at expiry.
[…]
    Debt: a broad class of fixed-income securities including variable return notes, medium-term notes, and government bonds, not otherwise classified as standard bonds.
[…]
    Equity Linked Note (ELN): a structured product whose returns are linked to the performance of an underlying equity or basket of equities.
[…]
    Other: a catch-all for instrument types not covered by the above classes, enabling schema extension and novelty detection.
Figure 10: Coverage vs. Utilization by Batch Size. The number of unmatched holdings covered (bars, left axis) decreases with increasing batch size, while the fraction of schema suggestions utilized (line, right axis) increases. This highlights a tradeoff: small batches are more exhaustive in coverage, but large batches are more efficient—yielding fewer “wasted” schema suggestions.
Batch Name Diversity Schema Diversity
Size Avg Min Max Avg Min Max
10 25.94 0 82 331.80 0 1387
50 22.67 0 78 313.41 0 1305
100 24.21 0 71 350.32 0 1569
250 22.60 0 55 342.97 0 1230
500 20.35 0 54 246.40 0 737
Table 8: Diversity metrics of unique schema suggestions for varying batch sizes. We report the average/minimum/maximum pairwise Levenshtein distance; “schema” metrics are over the entire generated schema, “name” is on the generated holding class name.
Appendix I Ablation Strategies
Raw Text Prompting.
For the baseline ablation, we prompt the LLM solely with the raw page text, asking whether a portfolio table is present via a simple yes/no detection prompt. Upon affirmative detection, the LLM is instructed to extract a portfolio table from the same text, returning the result as a structured object with a portfolio field, but without access to any schema or structural guidance. This strategy measures the LLM’s extraction performance in the absence of schema scaffolding or explicit reasoning.
Structured Chain-of-Thought (CoT).
To assess the impact of explicit reasoning on table detection, we prompt the LLM with the page text and require a structured Pydantic output containing both a chain-of-thought explanation (table_chain_of_thought) and a boolean indicating the presence of a portfolio table (has_portfolio_table). This ablation isolates the effect of minimal schema guidance and encourages the model to make its decision transparent through explicit intermediate reasoning. Upon positive detection, extraction is performed identically to the baseline, without additional schema context.
Full Schema Prompting.
In this ablation, we inject the complete Portfolio Pydantic schema directly into the detection prompt, alongside the page text. The LLM is instructed to reason about the presence of a portfolio table, outputting a chain-of-thought (chain_of_thought), a boolean detection (has_portfolio_table), and, if present, an extracted portfolio object conforming to the provided schema. This strategy evaluates the effect of strong schema supervision on both detection and extraction performance, requiring the model to both reason and map raw text into the structured schema within a single step.
Direct Schema Application.
For the final ablation, we bypass explicit table detection and directly apply the Portfolio schema extraction to every page. The LLM is prompted to extract a portfolio table from the provided text and return a Pydantic object with a portfolio field, irrespective of any prior detection or reasoning. Extraction is considered successful if any portion of the schema can be instantiated from the text. This approach evaluates schema-constrained extraction in the absence of explicit detection or intermediate supervision.
Appendix J Aggregation and Conflict Resolution of Schema Suggestions
After the LLM returns a batch of schema suggestions, we aggregate and cluster similar proposals as follows:
  1. 1.
    Deduplication: Suggestions with Levenshtein similarity $\geq 0.9$ (on class name and field structure) are merged.
  2. 2.
    Clustering: All proposals are clustered by semantic similarity of class names and required fields, using LLMs as the decision process.
  3. 3.
    Selection: For each cluster, the most frequent or most comprehensive schema suggestion is selected.
  4. 4.
    Validation: Each selected schema is validated by re-extracting unmatched holdings; suggestions that do not match any holding are dropped.
  5. 5.
    Manual review: If ambiguity remains, a manual review is triggered for final decision. We validated 64 resolved schemas for the second phase of extraction. Listing 15 displays the reconciled JSON schema for Forward Currency Contract, corresponding Pydantic model via pydantic.create_model, and several re-extracted holdings.
Appendix K Schema Suggestion Diversity
Table 8 summarizes the diversity among schema suggestions across batch sizes. Moderate batch sizes (100–250) achieve the highest average and maximum diversity, while the largest batch size (500) yields the lowest. This indicates that extremely large batches tend to generate more homogeneous or redundant suggestions, while moderate batches foster a broader range of candidate schemas.
Figure 11: Left: Name diversity (average, minimum, and maximum pairwise Levenshtein distance) among schema suggestions for varying batch sizes. Right: Schema diversity for the same. Moderate batch sizes (100–250) maximize diversity, while very large batches yield more homogeneous outputs.
Appendix L Example Holdings Tables
We show example holdings tables, alongside TASER’s extractions in Figures 19 - 37.
Batch Coverage (Holdings) Utilization (Schemas) Reported Collision
Size Count % Covered NAV (%) Total # # Utilized % Collisions Rate (%)
10 16,942 96.1 99.65 867 251 29.0 2,409 73.5
50 7,416 42.1 35.35 586 240 41.0 442 57.0
100 7,311 41.5 35.39 495 217 43.8 218 30.6
250 6,955 39.5 35.46 351 156 44.4 75 17.6
500 6,349 36.0 35.10 184 109 59.2 30 14.0
Table 9: Schema Utilization Efficiency. We report the proportion of generated schema suggestions that were utilized (i.e., matched at least one holding), for matching. Larger batch sizes result in a higher fraction of utilized schemas, suggesting that bulkier suggestion rounds are more efficient at targeting actionable schemas, albeit at the expense of overall diversity and coverage.
Detection Extraction End to End
Method Tokens Latency (s) Tokens Latency (s) Tokens Latency (s)
TASER
(a) Raw Text Prompting 1,495 0.33 5,414 20.37 6,909 20.69
(b) Structured CoT 1,514 1.70 5,440 20.20 6,954 21.90
(c) Full Schema Prompting 5,706 1.58 5,235 20.48 10,941 22.07
(d) Direct Schema Application 5,693 21.47 5,693 21.47
Table 10: Efficiency comparison of each ablation strategy. We report the token consumption and inference latency for detection, extraction, and end-to-end processing. Raw Text Prompting minimizes detection cost (1,495 tokens, 0.33 s) and achieves a total pipeline latency of 20.69 s; Structured CoT incurs additional reasoning overhead (1.70 s) with similar extraction performance; Full Schema Prompting uses the most detection tokens (5,706) but maintains comparable end-to-end latency (22.07 s); Direct Schema Application skips the detection stage entirely, applying schema validation directly in extraction. Dashes (—) indicate stages not performed by the method.
1# Ablation 1: Raw Text Prompting
2detection_prompt = (
[…]
5)
6
7
8# TableDetectionResponse Pydantic Model
9class TableDetectionResponse(BaseModel):
10 table_chain_of_thought: str = Field(...,
11 description="Chain of thoughts on if the page text contains table-like content")
12 has_portfolio_table: bool = Field(...,
13 description="True if the page has a holdings table, False otherwise")
14
15
16# Ablation 2: Structured Chain-of-Thought (CoT)
17detection_prompt = (
18 "Analyze the following text and determine if it contains a portfolio table. "
19 "Provide your chain of thought and final decision in a structured output "
[…]
22)
23
24
25# Ablation 3: Full Schema Prompting
26detection_prompt = (
27 "Using the provided Portfolio JSON schema, analyze the following text and "
28 "if it can be extracted into that schema. Provide your chain of thought. "
29 "You will output a response model object including ’chain_of_thought’, "
[…]
33)
34
35
36# Ablation 4: Direct Schema Application
37detection_prompt = (
38 "Extract a portfolio table from the following text following the Portfolio schema. "
[…]
41)
Figure 12: Detection prompts for all ablation strategies. Each section is labeled with its corresponding ablation strategy
Instrument Category Count Example
Equities 28,737 Taiwan Semiconductor Manufacturing
Debt 17,105 US Treasury 4.69% 09/05/2024
Unmatched Instruments 16,822 EUR
Forwards 8,023 Bought USD Sold KRW at 0.00072513
Options 977 Written Call Unilever 4050
Futures 720 US 5 Year Bond Future
Swaps 776 Pay fixed 3.026% receive float. (1d SOFR)
ELNs 292 BNP (Laobaixing Pharm. Chain (A)) ELN 22/07/2024
Table 11: Distribution of instrument categories in the dataset, with an example for each.
Listing 1: Portfolio schema with all matched instrument types.
1from enum import Enum
2from typing import Optional, List, Literal
3from pydantic import BaseModel, Field
4from datetime import datetime
5
6class BaseInstrument(BaseModel):
7 cusip: Optional[str] = Field(None, description="CUSIP identifier")
8 isin: Optional[str] = Field(None, description="International Securities Identification Number")
9 ticker: Optional[str] = Field(None, description="Ticker Symbol")
10 description: Optional[str] = Field(None, description="Description or name of the instrument")
11 quantity: Optional[float] = Field(None, description="Number of units held")
12 market_value: Optional[float] = Field(None, description="Market value of the holding")
13
14class Equity(BaseInstrument):
15 instrument_type: Literal["Equity"] = "Equity"
16 exchange: Optional[str] = Field(None, description="Trading exchange for the equity")
17
18class Option(BaseInstrument):
19 instrument_type: Literal["Option"] = "Option"
20 underlying: Optional[str] = Field(None, description="Identifier for the underlying asset")
21 strike_price: Optional[float] = Field(None, description="Strike price of the option")
22 expiration_date: Optional[datetime] = Field(None, description="Expiration date of the option")
23 option_type: Optional[str] = Field(None, description="Call or Put option")
24
25class Swap(BaseInstrument):
26 instrument_type: Literal["Swap"] = "Swap"
27 notional_amount: Optional[float] = Field(None, description="Notional amount of the swap")
28 fixed_rate: Optional[float] = Field(None, description="Fixed rate component (if applicable)")
29 floating_rate_index: Optional[str] = Field(None, description="Index used for floating rate leg")
30 maturity_date: Optional[datetime] = Field(None, description="Maturity date of the swap")
31 counterparty: Optional[str] = Field(None, description="The name of the counterparty")
32
33class Forward(BaseInstrument):
34 instrument_type: Literal["Forward"] = "Forward"
35 forward_price: Optional[float] = Field(None, description="Agreed forward price")
36 settlement_date: Optional[datetime] = Field(None, description="Settlement date for the forward")
37
38class Future(BaseInstrument):
39 instrument_type: Literal["Future"] = "Future"
40 contract_size: Optional[int] = Field(None, description="Size of the contract")
41 expiration_date: Optional[datetime] = Field(None, description="Expiration date of the future")
42
43class Debt(BaseInstrument):
44 instrument_type: Literal["Debt"] = "Debt"
45 coupon_rate: Optional[float] = Field(None, description="Annual coupon rate of the debt/bond")
46 maturity_date: Optional[datetime] = Field(None, description="Maturity date of the debt/bond")
47 issuer: Optional[str] = Field(None, description="Issuer of the debt/bond")
48
49class EquityLinkedNote(BaseInstrument):
50 instrument_type: Literal["Equity Linked Note"] = "Equity Linked Note"
51 issuer: Optional[str] = Field(None, description="Issuer of the ELN")
52 product: Optional[str] = Field(None, description="Underlying product of the ELN")
53 maturity_date: Optional[datetime] = Field(None, description="Maturity date of the ELN")
Listing 2: Main Portfolio Model with Unmatched (Other) Holdings class
1class Other(BaseModel):
2 description: str = Field(..., description="Text of the unknown instrument.")
3 name: str = Field(..., description="Suggested classification of the description or type")
4 market_value: Optional[float] = Field(None, description="Market value associated with the instrument")
5
6class Portfolio(BaseModel):
7 fund_name: Optional[str] = Field(None, description="Name of the fund that the portfolio belongs to")
8 value_in_thousands: bool = Field(False, description="True if the market value is based on thousands")
9 equities: Optional[List[Equity]] = Field(default_factory=list, description="List of equities")
10 options: Optional[List[Option]] = Field(default_factory=list, description="List of options")
11 swaps: Optional[List[Swap]] = Field(default_factory=list, description="List of swaps")
12 forwards: Optional[List[Forward]] = Field(default_factory=list, description="List of forwards")
13 futures: Optional[List[Future]] = Field(default_factory=list, description="List of futures")
14 debt: Optional[List[Debt]] = Field(default_factory=list, description="List of debt instruments")
15 elns: Optional[List[EquityLinkedNote]] = Field(default_factory=list, description="List of equity linked notes")
16 other_instruments: Optional[List[Other]] = Field(default_factory=list, description="The list of instruments that do not match any other type")
Figure 13: Class diagram of the initial Portfolio schema, showing the top‑level Portfolio containing a collection of Instrument objects, each subclassed into specific security types (Equity, Bond, Future, Forward, Swap, Option) to capture their unique attributes.
1# Prompt template for Recommender Agent, using batch size parameterization
2
3def recommender_agent_prompt(
4 portfolio_schema: dict,
5 unmatched_holdings: list,
6 batch_size: int,
7 start: int = 0,
8 previous_suggestions: list = None
9):
10 return """
11You are a schema refinement assistant for financial tables. Your task is:
12- Review a batch of {batch_size} unmatched financial holdings.
13- Given the current schema (JSON below), propose new classes or modifications so each holding can be classified.
14- If a holding matches a previously suggested class, propose new optional fields if needed.
15- Return your schema suggestions as a list of Pydantic SchemaSuggestion model objects.
16
17Current Portfolio Schema:
18{Portfolio.model_json_schema()}
19
20Batch of unmatched holdings:
21{unmatched_holdings[start : start + batch_size]}
22
23Previously seen suggestions (optional, from prior batches):
24{previous_suggestions if previous_suggestions else None}
25
26For each unique holding, propose:
27- A new schema class, or a modification to an existing class (add or refine fields).
28- Specify all required and optional fields with Python type hints.
29- If similar to an earlier suggestion, mark only new fields as optional.
30- Provide a sample match (the original holding string).
31- Output format: a Python list of SchemaSuggestion objects, as defined below.
32"""
33
34
35class SchemaSuggestion(BaseModel):
36 name: str # Name of new or modified schema class
37 suggested_schema: str # JSON schema for the instrument.
38 example: str # Example instrument seen in unmatched holdings
Figure 14: Recommender Agent schema suggestion prompt, output model, and example LLM response. The agent sees a batched portion of unmatched holdings to recommend new alterations to the Portfolio schema. This prompt is batch-specific and may include previous_suggestions for cross-batch refinement and de-duplication.
Listing 3: Currency Forward Generated JSON Schema
1{
2 "title": "Currency Forward",
3 "type": "object",
4 "properties": {
5 "description": {
6 "type": "string",
7 "title": "Description",
8 "description": "Description or name of the currency forward"
9 },
10 "market_value": {
11 "anyOf": [
12 { "type": "number" },
13 { "type": "null" }
14 ],
15 "title": "Market Value",
16 "description": "Market value of the currency forward",
17 "default": null
18 },
19 "instrument_type": {
20 "type": "string",
21 "title": "Instrument Type",
22 "const": "Currency Forward",
23 "default": "Currency Forward"
24 },
25 "currency_pair": {
26 "anyOf": [
27 { "type": "string" },
28 { "type": "null" }
29 ],
30 "title": "Currency Pair",
31 "description": "Currency pair involved in the forward contract",
32 "default": null
33 },
34 "forward_rate": {
35 "anyOf": [
36 { "type": "number" },
37 { "type": "null" }
38 ],
39 "title": "Forward Rate",
40 "description": "Agreed forward rate",
41 "default": null
42 },
43 "settlement_date": {
44 "anyOf": [
45 { "type": "string", "format": "date-time" },
46 { "type": "null" }
47 ],
48 "title": "Settlement Date",
49 "description": "Settlement date for the currency forward",
50 "default": null
51 }
52 }
53}
Listing 4: Currency Forward Pydantic Model
1class CurrencyForward(BaseModel):
2 description: str
3 market_value: Optional[float]
4 instrument_type: str = "Currency Forward"
5 currency_pair: Optional[str]
6 forward_rate: Optional[float]
7 settlement_date: Optional[datetime]
Listing 5: Refined Extraction
1# Raw inputs
2"Bought EUR Sold USD at 0.93035372 11/06/2024"
3"Bought USD Sold GBP at 1.25473636 31/05/2024"
4"Bought GBP Sold USD at 0.79368122 16/05/2024"
5
6# Extracted as fields
7{
8 "description": "Bought EUR Sold USD at 0.93035372 11/06/2024",
9 "market_value": -282515.0,
10 "instrument_type": "Currency Forward",
11 "currency_pair": "EUR/USD",
12 "forward_rate": 0.93035372,
13 "settlement_date": "2024-06-11T00:00:00"
14},
15{
16 "description": "Bought USD Sold GBP at 1.25473636 31/05/2024",
17 "market_value": 20651.0,
18 "instrument_type": "Currency Forward",
19 "currency_pair": "USD/GBP",
20 "forward_rate": 1.25473636,
21 "settlement_date": "2024-05-31T00:00:00"
22},
23{
24 "description": "Bought GBP Sold USD at 0.79368122 16/05/2024",
25 "market_value": 1429313.0,
26 "instrument_type": "Currency Forward",
27 "currency_pair": "GBP/USD",
28 "forward_rate": 0.79368122,
29 "settlement_date": "2024-05-16T00:00:00"
30}
Figure 15: Left: Final Currency Forward JSON schema. Top right: Equivalent Pydantic model. Bottom right: Example input string and its extraction into schema fields. This demonstrates schema-driven parsing of text into structured portfolio data. A currency forward contract is a financial instrument in the foreign exchange market that locks in the price at which an entity can buy or sell a currency at a future date.
Figure 16: Holdings Table Example 1
Market Coupon Maturity
Description Quantity Value Type Rate Date Issuer Debt Type
Poland Government 0.25% 25/10/2026 329000 58000 Debt 0.25 10/25/2026 Poland Government Government Bond
Poland Government 1.25% 25/10/2030 806000 127000 Debt 1.25 10/25/2030 Poland Government Government Bond
Poland Government 1.75% 25/04/2032 822000 127000 Debt 1.75 04/25/2032 Poland Government Government Bond
Poland Government 2.25% 25/10/2024 1688000 329000 Debt 2.25 10/25/2024 Poland Government Government Bond
Poland Government 5.75% 25/04/2029 499000 103000 Debt 5.75 04/25/2029 Poland Government Government Bond
Romania Government 3.25% 24/06/2026 1035000 168000 Debt 3.25 06/24/2026 Romania Government Government Bond
Romania Government 3.65% 28/07/2025 175000 30000 Debt 3.65 07/28/2025 Romania Government Government Bond
Romania Government 3.65% 24/09/2031 610000 89000 Debt 3.65 09/24/2031 Romania Government Government Bond
Romania Government 4.75% 24/02/2025 590000 101000 Debt 4.75 02/24/2025 Romania Government Government Bond
Romania Government 6.7% 25/02/2032 415000 74000 Debt 6.7 02/25/2032 Romania Government Government Bond
Republic of South Africa 8% 31/01/2030 3362492 133000 Debt 8 01/31/2030 Republic of South Africa Government Bond
Republic of South Africa 8.5% 31/01/2037 6431353 216000 Debt 8.5 01/31/2037 Republic of South Africa Government Bond
Republic of South Africa 8.75% 31/01/2044 3011713 96000 Debt 8.75 01/31/2044 Republic of South Africa Government Bond
Republic of South Africa 8.75% 28/02/2048 3687306 116000 Debt 8.75 02/28/2048 Republic of South Africa Government Bond
Republic of South Africa 9% 31/01/2040 3143167 105000 Debt 9 01/31/2040 Republic of South Africa Government Bond
Thailand Government 2% 17/06/2042 3761000 74000 Debt 2 06/17/2042 Thailand Government Government Bond
Thailand Government 2.125% 17/12/2026 15929000 363000 Debt 2.125 12/17/2026 Thailand Government Government Bond
Thailand Government 3.3% 17/06/2038 3760000 90000 Debt 3.3 06/17/2038 Thailand Government Government Bond
Thailand Government 3.4% 17/06/2036 3668000 89000 Debt 3.4 06/17/2036 Thailand Government Government Bond
Thailand Government 3.775% 25/06/2032 4334000 108000 Debt 3.775 06/25/2032 Thailand Government Government Bond
Thailand Government 4.85% 17/06/2061 836000 23000 Debt 4.85 06/17/2061 Thailand Government Government Bond
UK Treasury 0% 08/01/2024 7000000 6994000 Debt 0 01/08/2024 UK Treasury Government Bond
UK Treasury 0% 22/01/2024 9500000 9473000 Debt 0 01/22/2024 UK Treasury Government Bond
UK Treasury 0% 19/02/2024 2900000 2880000 Debt 0 02/19/2024 UK Treasury Government Bond
UK Treasury 0% 29/04/2024 8200000 8062000 Debt 0 04/29/2024 UK Treasury Government Bond
US Treasury 4.125% 31/07/2028 13507300 10705000 Debt 4.125 07/31/2028 US Treasury Government Bond
US Treasury 4.5% 15/11/2033 19261400 15862000 Debt 4.5 11/15/2033 US Treasury Government Bond
Figure 17: Debt Extracted
Market Expiration
Description Quantity Value Type Date
CBT US 10 Year Ultra Future March 2024 181 236000 Future 03/01/2024
CBT US Ultra Bond (CBT) March 2024+ -28 0 Future 03/01/2024
EUX DAX Index Future March 2024 8 -23000 Future 03/01/2024
ICF Long Gilt Future March 2024 -21 -4000 Future 03/01/2024
NYF Mini MSCI Emerging Market Future March 2024 100 130000 Future 03/01/2024
Figure 18: Futures Extracted
Market Strike Expiration Option
Description Quantity Value Type Price Date Type
S&P 500 Index Put Option 4250 February 2024 -27 -13000 Option 4250 02/01/2024 Put
S&P 500 Index Put Option 4350 March 2024 -27 -40000 Option 4350 03/01/2024 Put
S&P 500 Index Put Option 4500 February 2024 27 36000 Option 4500 02/01/2024 Put
S&P 500 Index Put Option 4600 March 2024 27 94000 Option 4600 03/01/2024 Put
Figure 19: Options Extracted
Figure 20: Holdings Table Example 2
Market
Description Quantity Value Type Exch.
Whitehaven Coal 210649 1045544 Equity AU
China Merchants Energy Shipping (A) 835101 952512 Equity CN
Transocean 186877 801702 Equity US
China Merchants Energy Shipping 221199 252299 Equity CN
Franco-Nevada (USA) 15736 1950792 Equity CA
Wheaton Precious Metals 31475 1914624 Equity CA
Baoshan Iron & Steel (A) 1049711 1033923 Equity CN
Hindalco Industries 100895 910253 Equity IN
Zijin Mining Group (H) 66000 148118 Equity CN
Agnico Eagle Mines (US) 1244 99756 Equity CA
Container Corp of India 70391 769624 Equity IN
Techtronic Industries 32500 487513 Equity HK
Taiwan Semiconductor Manufacturing 68000 2053064 Equity TW
NAVER 15346 1976673 Equity KR
Galaxy Entertainment Group 408000 2027380 Equity HK
Yum China Holdings 37955 1729989 Equity CN
Focus Media Information Technology (A) 1342200 1342677 Equity CN
ANTA Sports Products 97200 1159888 Equity CN
Focus Media Information Technology 1094891 1095280 Equity CN
Sea ADR 9468 887814 Equity SG
Yum China Holdings 5300 248116 Equity CN
Kweichow Moutai (A) 9100 2238717 Equity CN
CP ALL (F) 739800 1501788 Equity TH
Universal Robina 437420 809466 Equity PH
Indofood CBP 165500 134750 Equity ID
Kweichow Moutai (A) 100 24601 Equity CN
Bangkok Dusit Medic Service (F) 836100 777695 Equity TH
HDFC Bank 100285 2064120 Equity IN
HDFC Bank ADR 28877 1812609 Equity IN
Hong Kong Exchanges & Clearing 29100 1199421 Equity HK
AIA Group 17000 149937 Equity HK
China Overseas Land & Investment 828000 1669841 Equity HK
Figure 21: Equities Extracted
Market Maturity
Description Quantity Value Type Issuer Product Date
CITI (Focus Media Inf. Tech (A)) ELN 12/02/2026 711200 711451 Equity Linked Note CITI Focus Media Inf. Tech (A) 12/02/2026
UBS (Focus Media Inf. Tech (A)) ELN 06/02/2025 368898 369028 Equity Linked Note UBS Focus Media Inf. Tech (A) 02/06/2025
HSBC (Focus Media Inf. Tech (A)) ELN 16/05/2025 157600 157656 Equity Linked Note HSBC Focus Media Inf. Tech (A) 05/16/2026
Figure 22: ELNs Extracted
Market Forward Settlement
Description Quantity Value Type Price Date
Bought EUR Sold USD at 0.90292802 16/10/2024 408560 2806 Forward 0.90292802 10/16/2024
Bought EUR Sold USD at 0.89751876 16/10/2024 18830 16 Forward 0.89751876 10/16/2024
Bought EUR Sold USD at 0.89544132 16/10/2024 13513 -20 Forward 0.89544132 10/16/2024
Bought EUR Sold USD at 0.89300230 16/10/2024 21836 -92 Forward 0.89300230 10/16/2024
Bought USD Sold EUR at 1.10609821 16/10/2024 12489 -101 Forward 1.10609821 10/16/2024
Figure 23: Forwards Extracted
Market Contract Expiration
Description Quantity Value Type Size Date
MSCI Malaysia Index Future 20/12/2024 1590630 58275 Future 1590630 12/20/2024
IFSC Nifty 50 Index Future 31/10/2024 -674245 4628 Future 674245 10/31/2024
MSCI Thailand Index Future 20/12/2024 117630 3600 Future 117630 12/20/2024
S&P500 Emini Index Future 20/12/2024 -1448000 -25350 Future 1448000 12/20/2024
Figure 24: Futures Extracted
Market Strike Expiration Option
Description Quantity Value Type Underlying Price Date Type Ticker
Purchased Put Nvidia 95 21/03/2025 35 19250 Option Nvidia 95 03/21/2025 Put Nvidia
Purchased Put Taiwan Semic Mfg ADR 155 20/12/2024 14 7896 Option Taiwan Semic Mfg ADR 155 12/20/2024 Put Taiwan Semic Mfg ADR
Written Call Tencent Holdings 450 30/10/2024 -16 -3171 Option Tencent Holdings 450 10/30/2024 Call Tencent Holdings
Written Call Alibaba Group Holding 110 30/10/2024 -13 -4526 Option Alibaba Group Holding 110 10/30/2024 Call Alibaba Group Holding
Written Call Techtronic Industries 115 30/10/2024 -14 -5359 Option Techtronic Industries 115 10/30/2024 Call Techtronic Industries
Written Call AIA Group 65 30/10/2024 -13 -9149 Option AIA Group 65 10/30/2024 Call AIA Group
Written Call AIA Group 62.5 30/10/2024 -13 -11841 Option AIA Group 62.5 10/30/2024 Call AIA Group
Written Call NVIDIA 125 21/03/2025 -35 -58100 Option NVIDIA 125 03/21/2025 Call NVIDIA
Figure 25: Options Extracted
Figure 26: Holdings Table Example 3
Market
Description Value Type
USD/CHF 0.1 million -1467.11 Forward
USD/DKK 0.4 million -627.32 Forward
USD/GBP 0.1 million -1408.9 Forward
USD/JPY 0.6 million -167.24 Forward
USD/NOK 1.6 million -6901 Forward
USD/SEK 0.2 million -505.44 Forward
Figure 27: Forwards Extracted
Market
Description Type Value
Cash at bank Other 1844776
Demand deposits at Depositary - EUR deposits Other 534181.49
Deposits in other EU/EEA currencies - Danish krone Other 74061.35
Deposits in other EU/EEA currencies - Norwegian krone Other 186487.92
Deposits in other EU/EEA currencies - Swedish krona Other 74121.02
Deposits in non-EU/EEA currencies - British pound Other 214117.1
Deposits in non-EU/EEA currencies - Hong Kong dollar Other 2061.23
Deposits in non-EU/EEA currencies - Japanese yen Other 2557.6
Deposits in non-EU/EEA currencies - Canadian dollar Other 81066.58
Deposits in non-EU/EEA currencies - Swiss franc Other 74577.41
Deposits in non-EU/EEA currencies - U.S. dollar Other 601544.3
Dividends/Distributions receivable Other 300071.97
Prepaid placement fee Other 879371.08
Receivables from exceeding the expense cap Other 22774.37
Other receivables Other 2326.23
Receivables from share certificate transactions Other 42598.6
Liabilities from cost items Other -566947.6
Liabilities from share certificate transactions Other -743025.51
Figure 28: Other Instruments Extracted
Figure 29: Holdings Table Example 4
Market
Description Value Type
Novo Nordisk 41930000 Equity
ASML 34798000 Equity
SAP 31184000 Equity
TotalEnergies 24622000 Equity
Siemens 23147000 Equity
UniCredit 23012000 Equity
Deutsche Boerse 19999000 Equity
Munich Re 19512000 Equity
Anheuser-Busch InBev 18885000 Equity
CRH 18877000 Equity
Sanofi 18857000 Equity
BNP Paribas 17815000 Equity
Schneider Electric 17192000 Equity
Novartis 17036000 Equity
Alcon 16833000 Equity
Safran 16448000 Equity
National Grid 16195000 Equity
Airbus 15578000 Equity
Compass 15493000 Equity
LVMH Moët Hennessy Louis Vuitton 15141000 Equity
Infineon 14905000 Equity
SGS 14327000 Equity
ASM International 13782000 Equity
Cellnex Telecom 12982000 Equity
Aena 12179000 Equity
KONE 12163000 Equity
Ryanair 11974000 Equity
Danone 11697000 Equity
Roche 11592000 Equity
Holcim 11387000 Equity
British Land 11252000 Equity
Smurfit Westrock 11087000 Equity
BAWAG 10607000 Equity
Adidas 10392000 Equity
Beiersdorf 10390000 Equity
VAT Group 10143000 Equity
Saint-Gobain 9829000 Equity
DSV 9267000 Equity
Syensqo 8991000 Equity
Anglo American 8085000 Equity
Hermès 7529000 Equity
Rheinmetall 7456000 Equity
Nestlé 7222000 Equity
Galderma 7028000 Equity
Bayer 6988000 Equity
Stellantis 5689000 Equity
Figure 30: Equities Extracted
Figure 31: Holdings Table Example 5
Description Quantity Market Value Type
Verbund 168302 11168000 Equity
Ageas 636980 24854000 Equity
Azelis Group 621052 13974000 Equity
Galapagos 139674 5397000 Equity
Umicore 526316 16565000 Equity
Novozymes B 335786 13970000 Equity
Fortum 2175101 28338000 Equity
Neste 495924 19882000 Equity
Outokumpu 4402518 19859000 Equity
Carrefour 981701 14413000 Equity
Danone 395995 17454000 Equity
Pernod Ricard 97294 16726000 Equity
Societe Generale 885047 19377000 Equity
Ubisoft Entertainment 666784 12290000 Equity
Worldline 692255 25562000 Equity
Bayer 829596 42218000 Equity
Beiersdorf 307918 29865000 Equity
Fresenius 661906 16613000 Equity
GEA Group 621022 22067000 Equity
Knorr-Bremse 321638 17138000 Equity
MTU Aero Engines 137059 27000000 Equity
Porsche Automobil Holding Preference 510872 24911000 Equity
Siemens Energy 1001433 16767000 Equity
Software 473266 10792000 Equity
Wacker Chemie 93580 11105000 Equity
Zalando 481447 18162000 Equity
Bank of Ireland Group 2251106 18185000 Equity
Intesa Sanpaolo 6737795 13635000 Equity
ASM International 38474 9953000 Equity
BE Semiconductor Industries 533800 30856000 Equity
CNH Industrial 261421 3714000 Equity
Koninklijke Philips 1316703 18517000 Equity
OCI 87808 2398000 Equity
QIAGEN 879188 36463000 Equity
Aker BP 438070 11056000 Equity
Mowi 1093259 15766000 Equity
Amadeus IT Group 266403 13491000 Equity
CaixaBank 6213097 21527000 Equity
AAK 1263617 18364000 Equity
Billerud 1817978 18498000 Equity
Elekta B 2077018 11364000 Equity
Munters Group 1521430 12313000 Equity
Mycronic 492329 7752000 Equity
Saab B 522982 16880000 Equity
SKF B 1870356 27343000 Equity
Svenska Handelsbanken A 2919741 25123000 Equity
Tele2 B 2424889 17979000 Equity
Viaplay Group B 550681 9641000 Equity
Cie Financiere Richemont 408985 49902000 Equity
Novartis 609263 45498000 Equity
Swiss Re 253533 20755000 Equity
Figure 32: Equities Extracted
Description Type Market Value
Net other assets Other 12016000
Figure 33: Other Instruments Extracted
Figure 34: Holdings Table Example 6
Market Coupon Maturity
Description Quantity Value Type Rate Date Issuer
CBT US 10 Year Ultra Future March 2024 181 236000 Future 4250 02/01/2024 Poland Government
CBT US Ultra Bond (CBT) March 2024+ -28 0 Future 4250 03/01/2024 Poland Government
EUX DAX Index Future March 2024 8 -23000 Future 4350 03/01/2024
ICF Long Gilt Future March 2024 -21 -4000 Future 4500 02/01/2024
NYF Mini MSCI Emerging Market Future March 2024 100 130000 Future 4600 03/01/2024
Figure 35: Debt Extracted
Market Expiration
Description Quantity Value Type Date
US 10 Years Note Short futures contracts -35 55508 Future 02/03/2025
US 5 Years Note Short futures contracts -17 11953 Future 31/03/2025
Figure 36: Futures Extracted
Market Type Settlement
Description Quantity Value Type Date
USD/CAD Forward Contract 1 53468 Forward 31/01/2025
USD/EUR Forward Contract 1 482874 Forward 31/01/2025
USD/GBP Forward Contract 1 43213 Forward 31/01/2025
Figure 37: Forwards Extracted