### v4

# TASER: Table Agents for Schema‑guided Extraction and Recommendation

Abstract
Real-world financial filings report critical information about an entity’s investment holdings, essential for assessing that entity’s risk, profitability, and relationship profile. Yet, these details are often buried in messy, multi-page, fragmented tables that are difficult to parse, hindering downstream QA and data normalization. Specifically, 99.4% of the tables in our financial table dataset lack bounding boxes, with the largest table spanning 44 pages. To address this, we present TASER (Table Agents for Schema-guided Extraction and Recommendation), a continuously learning, agentic table extraction system that converts highly unstructured, multi-page, heterogeneous tables into normalized, schema-conforming outputs. Guided by an initial portfolio schema, TASER executes table detection, classification, extraction, and recommendations in a single pipeline. Our Recommender Agent reviews unmatched outputs and proposes schema revisions, enabling TASER to outperform vision-based table detection models such as Table Transformer by 10.1%. Within this continuous learning process, larger batch sizes yield a 104.3% increase in useful schema recommendations and a 9.8% increase in total extractions. To train TASER, we manually labeled 22,584 pages and 3,213 tables covering \$731.7 billion in holdings, culminating in TASERTab to facilitate research on real-world financial tables and structured outputs. Our results highlight the promise of continuously learning agents for robust extractions from complex tabular data.
TASER: Table Agents for Schema‑guided Extraction and Recommendation
Nicole Cho    Kirsty Fielding    William Watson Sumitra Ganesh   Manuela Veloso J.P. Morgan AI Research nicole.cho@jpmorgan.com
1 Introduction
Figure 1: Complexity of Holdings Table in Regulatory Filings. In the original format, multiple data attributes are displayed in a single line, with no bounding boxes, rendering the generation of structured outputs highly challenging. TASER enables the generation of structured outputs from highly variable, multi-page financial tables for complex instrument holdings. Negative quantities or market values denote short positions. See Appendix K for additional outputs.
Figure 2: Variety and complexity of financial tables. From leftmost column - for a single financial filing type, such as annual reports, there is no consistency among reports. Within each report, there are numerous table types with each type housing very different types of information. Even within a single table type (such as the Financial Holdings Table), there are numerous layout structures, as seen in the rightmost column. Due to the extreme heterogeneity of formatting, document layout, and table structure, traditional table extraction methods fail to perform for financial filings.
Financial documents, particularly annual regulatory filings for funds, house tables that govern \$68.9 trillion of investments globally (Investment Company Institute, 2024). By comparison, \$68.9 trillion is more than twice the total Gross Domestic Product (GDP) of the United States (\$29.1 trillion) (WorldBank, 2025). This critical data is housed in the Financial Holdings Table (Figure 1), which outlines the entirety of an entity’s investment holdings (U.S. Congress, 1934; EU Commission, 2019); this table has the highest row count (maximum 426 rows)–more than double the average row count of all other table types (Table 7). These Financial Holdings Tables are long and highly heterogeneous in layout (Figure 2). While generating structured outputs from these tables is critical for many regulatory and financial institutions to undertake basic QA (Question-Answering) tasks using an LLM (Large Language Model) or libraries such as pandas (Cho et al., 2024), there is a relative dearth of studies that focus on continual learning to extract from Financial Holdings Tables, compared to web or SQL tables (Herzig et al., 2020; Pasupat and Liang, 2015; Zhong et al., 2017). Therefore, the following challenges exist in terms of parsing Financial Holdings Tables into structured, machine-readable outputs: (1) One-to-many relationships between a document and the tables it houses exacerbate standard model performance for table detection or structure recognition tasks. (2) Financial Holdings Tables span across multiple pages, rendering models that operate at the page level inefficient. (3) Financial instruments are highly complex with nested hierarchies. Therefore, details are often clumped in a single cell as seen in Figure 1. (4) Tabular layouts are heterogeneous with no bounding boxes, mixing tables, text blocks, footnotes, and images, often without consistent labeling or alignment. 99.4% of tables in our dataset, TASERTab, lack bounding boxes to efficiently identify a single cell. These challenges motivate our agentic table extraction methodology capable of goal-driven parsing and self-refinement, continuously learning and reasoning from errors.
Contribution 1: We propose a continuously learning, agentic table extraction methodology, TASER (Table Agents for Schema-guided Extraction and Recommendation) that performs detection, classification, extraction, and recommendation in a single pipeline by leveraging the schema invoked as a tool call. TASER is layout-agnostic and can operate for tables of any format. We compare our methodology against predominant methodologies and report TASER’s 10.1% improvement over Table Transformer (Yang et al., 2022) for detection.
Contribution 2: We demonstrate the effectiveness of our Recommender Agent to continuously improve the initial schema - reflecting a tunable and continuous self-learning loop. Throughout our training, we found that small batches are optimal for providing diverse and comprehensive recommendations to the original schema–however, at the cost of redundant recommendations. In contrast, large batches drive high precision recommendations at the cost of diversity. Thus, our results establish that self-learning via agents for table extraction is tunable; through adjusting batch size, we can control schema refinement to maximize actionable coverage while minimizing redundancy.
Contribution 3: We have constructed a manually labeled dataset TASERTab of ground truth labels for 3,213 real-world Financial Holdings Tables amounting to \$731.7B in value. We sourced the filings from fund websites, labeled the total net assets for each fund, and recorded the span of each Financial Holdings Table. We believe that this is the first dataset of its kind to provide access to real-world financial tables side by side with structured outputs.
2 Related Work
Information & Table Extraction: Early information extraction relied on statistical models (HMMs (Borkar et al., 2001), CRFs (Lafferty et al., 2001), heuristics (Press, 2003), and graph-based layouts (Liu et al., 2019; Qian et al., 2019; Meuschke et al., 2023), but still struggle with complex, heterogeneous tables.
Table Representation Learning: Transformer-based table understanding and QA include TaPaS (Herzig et al., 2020), TaBERT (Yin et al., 2020), TaPEX (Liu et al., 2022), TURL (Deng et al., 2020), TUTA (Wang et al., 2021), and TableFormer (Yang et al., 2022). These methods encode text, structure, and layout, but few are benchmarked on long, dense, multi-page financial reports.
LLMs for Structured Data: General LLMs have strong performance for schema-conformant extraction via fine-tuning & prompting (Brown et al., 2020; Liu and Contributors, 2024), while multimodal approaches (LayoutLM (Xu et al., 2020, 2021; Huang et al., 2022), DONUT (Kim et al., 2022), DocFormer (Appalaraju et al., 2021), UniTable (Peng et al., 2024), and Table Transformer (Smock et al., 2021; Carion et al., 2020) improve layout awareness but still lag on long, fragmented tables (Zhao et al., 2024).
Financial Document Parsing: (Watson and Liu, 2020) has focused on table extraction from images while (Cho et al., 2024) has focused on expert agent pipelines. Large-scale benchmarks such as DocILE (Šimsa et al., 2023), BuDDIE (Wang et al., 2025) have also focused on financial documents.
Agentic and Recursive Extraction: Recent methods cast LLMs as agents capable of iterative extraction and self-correction (Shen et al., 2023; Roucher et al., 2025; Watson et al., 2023; Yuan and Xie, 2025). Prompt-based feedback, introspective refinement, and episodic memory frameworks (Madaan et al., 2023; Shinn et al., 2023; Yao et al., 2023) drive improvements in reasoning for complex extraction.
3 Methodology
3.1 System Architecture
Our methodology is composed of three core Large Language Model (LLM) agents, each with a distinct role. We conduct rigorous ablations to evaluate the importance of each agent.
  1. 1.
    Detector Agent: Identifies candidate pages containing Financial Holdings Tables leveraging the initial schema provided. The prompt is tuned to maximize recall to avoid missing any Financial Holdings Tables. We provide our prompts in the Appendix (Figure 12).
  2. 2.
    Extractor Agent: Processes detected pages by prompting the LLM with the current Portfolio schema embedded in the prompt context. The LLM’s output is validated inline against the schema using Pydantic & Instructor, producing a set of structured, type-checked instrument entries (Figure 1).
  3. 3.
    Recommender Agent: Reviews unmatched extractions containing both false and true positives. A false positive is a spurious extraction (e.g., headers/subtotals/footnotes/OCR noise or cells from non-holdings tables) that fails schema/type/consistency checks; a true positive is a valid holdings field from a genuine row that the current schema cannot yet classify but passes those checks. The agent first filters false positives by re-validating each candidate under the current schema; it proposes schema modifications for the remaining true positives and, per class, recommends the minimal change needed.
All agents interact through explicit artifacts: structured outputs, episodic error stacks, and schema definitions. Output validation is integrated into each agent’s forward pass via Instructor (Liu and Contributors, 2024). TASER implements a recursive feedback loop, where errors and unmatched holdings identified in the initial extraction are escalated to the Recommender Agent, which provides recommendations to refine the schema and triggers re-extraction. This loop repeats until all entries are matched. A schematic of the full agentic pipeline is shown in Figure 3.
Figure 3: TASER deployment architecture. Users submit extraction requests through a UI or API hosted in an ECS container, which enqueues an extraction task to SQS. An AWS Lambda function orchestrates the Ingest, Detection, Extraction, and Recommender Agents, persisting task status and schema cache in DynamoDB and writing intermediary and final outputs to S3; failed tasks are routed to a dead-letter queue (DLQ) for later inspection. The Recommender Agent analyzes the intermediary output file and associated error stack to propose schema enhancement recommendations. Users can accept or reject these recommendations; accepted updates are written back to the schema cache and automatically retrigger the extraction pipeline, enabling TASER to continuously refine its extraction artifacts over time.
3.2 Initial Schema Definition and Application
TASER’s extraction process is anchored by an explicit, user-modifiable Portfolio schema that defines the target structure for Holdings Tables. We implement this schema using Pydantic models; our initial schema reflects is informed by leveraging external knowledge (U.S. Congress, 1934). Each schema consists of a base Instrument model, subclassed for common asset types (e.g., Equity, Bond, Option, Swap, Forward, Future, Debt, and an Other class for uncategorized rows. Each subclass specifies instrument-specific fields and validation logic (see App. G).
Schema-Guided Extraction: For each candidate page, the Extractor Agent prompts the LLM with the current schema embedded in the prompt context. The LLM is instructed to return a structured output, which is immediately parsed and validated against the schema using Pydantic’s type checking and validation logic. Outputs that fail schema validation (e.g., missing fields, type errors, or undeclared instruments) are flagged.
Schema Recommendations for Iterative Refinement: We formalize schema refinement as an iterative, LLM-driven clustering process that updates the schema to accommodate unmatched or novel holdings discovered during extraction (Novikov et al., 2025; Zhang et al., 2023). At each iteration, the agent operates on the episodic error stack to propose schema modifications, and extraction is retried using the updated schema. This process continues until all entries are matched or no further improvements are possible. Let $H=\{h_{1},h_{2},\ldots,h_{N}\}$ denote the set of unmatched holdings, and let $\Sigma^{(0)}$ be the initial schema. For each iteration $\ell$:
[…]
    $H^{(\ell)}$: Unmatched holdings at iteration $\ell$.
[…]
    $\Sigma^{(\ell)}$: Current schema.
[…]
    $g_{\theta}$: LLM-based schema suggestion function.
[…]
    $B$: Batch size for error grouping.
The refinement loop (Algo. 1) proceeds as follows:
  1. 1.
    Partition $H^{(\ell)}$ into batches of size at most $B$.
  2. 2.
    For each batch, invoke $g_{\theta}$ with batch errors and $\Sigma^{(\ell)}$ to propose schema modifications.
  3. 3.
    Aggregate, cluster, and select recommendations
  4. 4.
    Update schema to $\Sigma^{(\ell+1)}$ and re-extract.
  5. 5.
    Update error stack and repeat until $H^{(\ell+1)}$ is empty or no new schema changes are suggested.
Figure 4: Schema-Guided Agentic Refinement Loop. The extraction pipeline begins with an Initial Schema Definition (v1), which guides the LLM Extractor Agent as it processes the raw Holdings Table to produce Declared Holdings. Holdings that do not match the schema are routed as Unmatched Holdings, triggering the generation of Schema Update Suggestions. These suggestions are reviewed, clustered, and aggregated by our Recommender Agent before updating the schema (v2), replacing the prior definition and closing the agentic feedback loop. This process enables continuous schema refinement and robust extraction.
Algorithm 1 LLM Iterative Schema Refinement
0: Unmatched holdings $H=\{h_{1},h_{2},\ldots,h_{N}\}$, initial schema $\Sigma^{(0)}$, LLM schema suggestion function $g_{\theta}$, batch size $B$, stopping criterion $T$
1: Initialize $\ell\leftarrow 0$
[…]
4:while not stopping criterion $T$ met do
5:  Partition $H^{(\ell)}$ into batches $H^{(\ell)}_{j}$ of size at most $B$
6:  $S^{(\ell)}\leftarrow\emptyset$ {Suggested schema modifications}
7:  for each batch $H^{(\ell)}_{j}$ do
8:   $S_{j}^{(\ell)}\leftarrow g_{\theta}(H^{(\ell)}_{j},\Sigma^{(\ell)})$
9:   $S^{(\ell)}\leftarrow S^{(\ell)}\cup S_{j}^{(\ell)}$
10:  end for
11:  $S_{\text{selected}}^{(\ell)}\leftarrow\text{AggregateAndSelect}(S^{(\ell)})$ {Aggregate suggestions}
12:  $\Sigma^{(\ell+1)}\leftarrow\text{UpdateSchema}(\Sigma^{(\ell)},S_{\text{selected}}^{(\ell)})$
[…]
14:  if $H^{(\ell+1)}=\emptyset$ then
15:   break
16:  end if
17:  $\ell\leftarrow\ell+1$
18:end while
19:return $\Sigma^{(\ell+1)}$
3.3 Ablation Strategies and Efficiency
We systematically ablate TASER to isolate the impact of schema-guided extraction, prompt engineering, and agentic feedback across four strategies:
  1. 1.
    Raw Text Prompting: The LLM is prompted only with the page text; extraction is based solely on a yes/no detection.
  2. 2.
    Structured Chain-of-Thought (CoT): Prompts include a minimal schema and few-shot examples, eliciting explicit reasoning traces before a final boolean decision.
  3. 3.
    Full Schema Prompting: The full Portfolio schema is embedded in the prompt, instructing the LLM to return structured, schema-conformant entries.
  4. 4.
    Direct Schema Application: The schema is directly applied to parsed page content without prior detection; extraction succeeds if any schema sub-model instantiates.
Table 2 reports detection and extraction via absolute dollar difference, and Table 10 compares computational efficiency in tokens and latency.
Provider Model Recall (%) Precision (%) F1 Score (%) Accuracy (%)
Camelot Hybrid 56.92 23.46 33.23 47.35
Microsoft Table Transformer 99.76 32.75 49.31 46.27
OpenAI gpt-4o-2024-11-20 100.00 43.43 59.44 66.35
OpenAI gpt-5-mini-2025-08-07 94.92 54.15 68.96 80.33
OpenAI gpt-4.1-2025-04-14 95.80 54.32 69.33 80.49
OpenAI gpt-5-nano-2025-08-07 95.97 55.63 70.44 81.46
OpenAI gpt-5-2025-08-07 95.97 68.16 79.71 88.75
Anthropic claude_sonnet-3-7 88.97 57.34 69.73 82.22
Amazon nova_pro-v1-0 85.90 69.45 76.84 88.07
Table 1: Detector performance across models on TASERTab. Recall, precision, F1, and accuracy are reported for baselines and the Detector Agent instantiated with different LLMs. gpt-4o-2024-11-20 attains perfect recall, while gpt-5-2025-08-07 achieves the best overall F1 and accuracy. nova_pro-v1-0 delivers the highest precision but at the cost of lower recall, illustrating the trade-off between missing holdings tables and avoiding false positives.
Method Recall (%) Precision (%) F1 (%) Accuracy (%) TAD (USD) Unaccounted
TASER (a) Raw Text Prompting 100.00 38.62 55.73 58.38 \$ 107,066,845 0.015%
(b) Structured CoT 100.00 34.42 51.21 50.10 \$ 120,577,458 0.016%
(c) Full Schema Prompting 100.00 43.43 59.44 66.35 \$ 102,836,797 0.014%
(d) Direct Schema Application 100.00 41.84 58.30 63.99 \$ 118,881,312 0.016%
Table 2: Detection and Extraction Performance Across Strategies. While all TASER ablations achieve perfect recall, Full Schema Prompting yields the highest precision (43.43%), F1 score (59.44%), and overall accuracy (66.35%), as well as the lowest total absolute difference (TAD) and unaccounted fraction, underscoring the value of embedding the complete Portfolio schema in the detection prompt. Percentage of unaccounted holdings is out of \$731.7 billion (ground truth). Lower TAD and unaccounted percentages indicate higher dollar‐value fidelity.
Figure 5: Left: Cumulative unique schemas per iteration; larger batches discover schemas rapidly but plateau quickly. Right: Cumulative unique schemas per unmatched holding seen; smaller batches ultimately yield more unique schemas but require more suggestions and generate more redundancy.
4 Experimental Setup
Detection Metrics: We report recall, precision, F1, and accuracy for table detection, prioritizing recall to avoid missing Financial Holdings Tables.
Extraction Metrics: We assess extraction completeness by comparing TASER’s outputs to ground truth labels. We manually label a total net asset value for each Holdings Table. We then compare this ground truth with our extractions, dubbed the total absolute difference (TAD).
Schema Refinement Metrics: Coverage is the fraction of unmatched holdings aligned with at least one schema suggestion, using RapidFuzz string similarity with a lenient ($\geq 70$) threshold. We also report the number of new matched holdings after re-extraction with the suggested schemas added to Portfolio. Diversity is the average pairwise Levenshtein distance between suggestion attributes (name and generated schema). Collision rate denotes the proportion of duplicate suggestions.
Dataset and Model: We curate a diverse corpus totaling 22,584 pages, 28M tokens, and \$731.7B in holdings. Among 3,213 tables, 57.53% exhibit hierarchical structure (via spanning cells). All Holdings Tables (100%) are hierarchical. While 39% of portfolios are single-page, 60.2% span multiple pages. The average length is 3.24 pages ($\sigma=3.41$, max = 19). This variability underscores the need for multi-page detection and consolidation. Unless explicitly stated otherwise, all experiments use gpt-4o-2024-11-20 as the LLM.
5 Results and Discussion
5.1 Quantitative Evaluation
Detection: Table 2 shows that all TASER ablations achieve perfect recall ($\sim$100%), but precision ranges from 32.8% (Table Transformer) up to 43.4% (Full Schema Prompting), driving F1 scores between 49.3% and 59.4%. Embedding the full Portfolio schema in the prompt boosts precision by over 10% relative to the vision‑only baseline and yields the highest F1 (59.4%) and accuracy (66.4%), demonstrating that in‑context schema guidance is critical.
Extraction: Table 2 confirms schema-anchored extraction improves dollar‑value fidelity. Full Schema Prompting attains the lowest absolute difference (\$102.8M) and smallest unaccounted share (0.014%), outperforming Raw Text Prompting (\$107.1M, 0.015%) and Structured CoT (\$120.6M, 0.016%). Direct Schema Application (skipping detection) incurs a higher error (\$118.9M; 0.016%) by parsing spurious non‑holding pages.
5.2 Success Highlights
Cross-Document Consistency: TASER classifies and extracts Holdings Tables despite varying titles (e.g., "Portfolio of Investments", "Schedule of Holdings", or "Investment Portfolio") and diverse structural formats. Despite the immense complexity of inputs, TASER consistently extracts and transforms these tables, ensuring that the final output appears as if sourced from a uniform set.
Contextual Understanding: TASER excels in handling contextual nuances, such as interpreting negative values denoted by parentheses (e.g., (140)) in zero-shot settings. Such domain-specific attributes are important for financial tables.
Extracting Intricate Semantics: TASER demonstrates a strong semantic understanding of financial terminology, which empowers it to extract accurately. For instance, TASER adeptly parsed the table entry “GBP 4,700,000 | UK Treasury 0% 19/02/2024 | 4668 | 1.48,” correctly identifying the holding as a bond and extracting its attributes: quantity, market value, coupon rate, maturity date, and issuer.
Figure 6: Reduction in TAD after resolving unmatched holdings. Remaining TAD is calculated after sequential reconciliation of unmatched holdings. Error reduction is achieved by resolving the most significant unmatched holdings (Figs. 8 & 9).
5.3 Batch Size Tradeoffs in Refinement
Figure 5 (left) reveals that larger batch sizes (250, 500) rapidly expand the schema. However, this early acceleration comes at the cost of early saturation, after which few new unique schemas are discovered. In contrast, smaller batches require more iterations to reach the same number of unmatched holdings seen, but continue yielding unique schemas, resulting in the highest diversity when normalized by data processed (Figure 5, right). This improvement in coverage, however, is offset by increased redundancy. As shown in Appendix Figure 7, smaller batches incur substantially more overlapping suggestions, reflecting a more granular and exploratory nature. Overall, these results highlight a key tradeoff: larger batches accelerate early discovery but plateau quickly, while smaller batches maximize cumulative schema diversity at the cost of redundancy and computation. Our results establish that schema refinement via agentic feedback is both tractable and tunable. This indicates that an adaptive batching strategy may be optimal: using larger batches to quickly identify high-yield schemas, followed by smaller batches for exhaustive diversity.
Schema Diversity and Utilization: Schema diversity, as measured by the average pairwise Levenshtein distance, is maximized for moderate batch sizes (100–250), as shown in Appendix J. While larger batch sizes (500) yield a higher proportion of utilized schemas—up to 59%—smaller, more diverse batches tend to have lower utilization rates (Table 9). Furthermore, the accretive gain in 402 additional unique schemas yielded only marginal improvements in holding coverage (6.1%). Figure 10 illustrates this tradeoff: smaller batch sizes cover more unmatched holdings at the expense of efficiency (96.1% coverage for 29.0% utilization at batch size 10), whereas larger batches achieve higher schema utilization (59.0% at batch size 500).
Improvements in TAD: Resolving the largest unmatched holdings yields a reduction in TAD of approximately 7–10% across batch sizes, with the majority of improvement achieved by reconciling just the top 10–20% of holdings (Table 5).
5.4 Deployment of TASER
We outline the deployment architecture of TASER in Figure 3, with additional system architecture details provided in Appendix A.
6 Conclusion
We present TASER for extracting complex Holdings Tables from documents through continual learning. Our high precision and recall across diverse layouts underscore the potential of agentic continual learning for financial table extraction.
Disclaimer
This paper was prepared for informational purposes by the Artificial Intelligence Research group of JPMorgan Chase & Co. and its affiliates ("JPMorgan”) and is not a product of the Research Department of JPMorgan. JPMorgan makes no representation and warranty whatsoever and disclaims all liability, for the completeness, accuracy or reliability of the information contained herein. This document is not intended as investment research or investment advice, or a recommendation, offer or solicitation for the purchase or sale of any security, financial instrument, financial product or service, or to be used in any way for evaluating the merits of participating in any transaction, and shall not constitute a solicitation under any jurisdiction or to any person, if such solicitation under such jurisdiction or to such person would be unlawful.
Limitations
Despite its strong performance, TASER remains susceptible to errors in low-resolution or scanned PDFs, where visual degradation can hinder accurate extraction. Ambiguities in financial documents, such as undefined asset classes or implicit references, pose challenges that cannot always be resolved without external knowledge or manual intervention. While recursive prompting enhances completeness, it introduces added latency and computational overhead. Additionally, TASER relies on prompt-based weak supervision due to the lack of fine-grained, labeled datasets for complex instrument types, which may limit generalization. Finally, TASER does not yet model interactions between table rows or instrument relationships beyond the schema level, which may affect downstream tasks such as portfolio risk analysis or exposure aggregation.
Appendix A TASER System Architecture
A.1 Overview
TASER is an event-driven microservice architecture composed of several semi-independent agents, such as text extraction and table detection modules. Each agent operates on a shared cloud infrastructure, which typically consists of an AWS Lambda function that listens to a queue (Amazon SQS), manages file operations in Amazon S3, and utilizes a dead-letter queue (DLQ) for retry handling.
A.2 User Interaction and Deployment
Users may interact with TASER either through a graphical user interface (UI) or directly via the application programming interface (API). Both the UI and API are deployed within an Amazon ECS cluster, with horizontal and vertical scaling enabled to respond to user demand. The UI is updated by periodically polling the API for task status, which is maintained in Amazon DynamoDB.
A.3 Task Cache
An additional component is the task cache, implemented using DynamoDB. The primary purpose of the cache is to reduce latency and operational costs by retrieving previously computed results. For all agents, the cached value is typically an S3 location indicating where the result of a prior execution is stored. The cache key varies by agent but generally includes model-specific information (e.g., model name, temperature settings) and the prompt used for generation. A cache entry for the table detection agent is structured as follows:
Key: {schema_id, prompt_id, document_text_id,
model_name, model_configs}
Value: S3 location
A.4 Design Rationale: API-Centric Orchestration
A core design decision for TASER was to utilize an API-centric approach for orchestrating workflows. An alternative would have been to configure each agent to automatically trigger an AWS Lambda function upon the arrival of a new file in S3 (either by directly configuring the S3 bucket or using AWS EventBridge). While this method is feasible, it means that any change to workflows would require reprovisioning AWS infrastructure. In contrast, with the current architecture, redesigning a workflow is as simple as refactoring the UI or API client code (for example, changing the order of API calls) and redeploying the application, without modifying the underlying cloud infrastructure.
Appendix B Camelot Table Parsing Modes
For completeness, we also compare TASER’s detection performance against the four table detection modes in Camelot [^1] https://github.com/camelot-dev/camelot. The best-performing variant (Hybrid) achieves an F1 score of 0.33, still below TASER’s weakest ablation (0.51). Full results are presented in Table 4. Note that our financial tables primarily consist of unruled, whitespace-separated tables with alignment-based structure. Below is a brief summary of each mode:
[…]
    Stream: Groups text using whitespace and y-axis alignment. Suitable for unruled tables, but yielded low precision on our data ($F1=21.6\%$).
[…]
    Lattice: Uses image-based line detection to extract ruled tables. Less effective for our dataset due to the rarity of bordered layouts ($F1=13.3\%$).
[…]
    Network: Detects tables via text alignment patterns using bounding boxes. Performs better on our format, which lacks explicit ruling ($F1=18.6\%$).
[…]
    Hybrid: Combines Network’s structure with Lattice’s grid refinement. Achieved the highest F1 score (33.23%) among Camelot modes, confirming the benefit of integrating both visual and alignment cues.
Figure 7: Cumulative collisions per unmatched holding; smaller batches incur more collisions, reflecting greater redundancy.
Appendix C TASER Annotation Process
We manually sourced each financial document directly from the fund entity’s public website, ensuring broad coverage across instrument types. Annotations were performed at the page, table, and holdings level (which may span hundreds of pages). For every filing and fund, we recorded the page-span for the portfolio of investments table and the net asset value across all holdings for that fund.
Appendix D TASER Dataset Release
TASER is built on public fund documents. Our release will include labels for the positions of holdings tables, the recorded net asset value, the fund name, multi-page spans, and a URL reference to the public fund document. Each pdf filing is hosted by the fund’s advisor, as required by regulation.
Batch Remaining TAD Reduction NAV
Size TAD (\$) (%) Extracted (\$)
500 94,843,638 7.8% 7,993,158
250 94,185,693 8.4% 8,651,103
100 95,985,588 6.7% 7,851,209
50 92,781,421 9.8% 10,025,376
10 93,032,549 9.6% 9,804,248
Table 5: Remaining Total Absolute Difference (TAD, \$) and Net Asset Value (NAV, \$) extracted from reconciled unmatched holdings by batch size.
Appendix E Document Preprocessing
For each PDF filing, TASER extracts raw text, layout metadata, and embedded images using a hybrid pipeline based on pdfplumber. Each page is parsed into normalized text blocks and layout primitives, preserving spatial relationships and read order. Minimal normalization is applied, including Unicode cleanup and header/footer removal. Each page object includes:
[…]
    Raw text blocks (reading order preserved)
[…]
    Bounding boxes and font metadata
[…]
    Embedded images (if any)
We apply Unicode normalization (NFKC), whitespace collapse, and filter out repeated headers/footers via regex matching. Optionally, OCR is performed if text extraction fails. Code and parameters are available upon request.
Figure 8: Heavy-tailed distribution of value recovery from unmatched holdings across batch sizes. We report the Lorenz curves for the cumulative fraction of value recovered as a function of the fraction of “other” holdings resolved. For all batch sizes, a small number of matches account for the vast majority of recovered net asset value, while most resolved holdings contribute negligibly. The bow of each curve away from the diagonal illustrates the extreme concentration of recoverable value in the “head,” characteristic of a heavy-tailed regime.
Figure 9: Cumulative recovery fraction vs. number of holdings resolved. Cumulative fraction of total value recovered as a function of the number of unmatched holdings resolved (log-log scale). The steep initial rise for each batch size indicates that the largest recoveries are concentrated among the first few resolved holdings; subsequently, improvement plateaus, indicating diminishing returns from resolving additional holdings in the long tail.
Appendix F Parallelization and Fund Construction
Extraction: To efficiently process large, multi-page filings, TASER employs parallelization (20 workers) at both the document and page levels. Each agent operates asynchronously across document batches: Detector and Extractor agents process candidate pages in parallel, while the Recommender agent operates downstream on the resulting artifacts.
Merging: For fund-level construction, extracted tables from consecutive pages are merged deterministically. Entity resolution is performed by matching predicted fund names and table headings across pages, while units and currencies are normalized to a consistent reporting standard through a boolean flag value_in_thousands. Partial extractions are reconciled using strict types in the response model, whose validation errors re-prompt the LLM on specific extraction errors to ensure a unified, schema-conformant portfolio representation for each fund.
Avg. # Tables Avg. Rows Avg. Columns Avg. Spanning
Dataset Per Topology Per Table Per Table Cells per Table
SciTSR 5.70 9.28 5.19 0.77
PubTabNet 4.13 14.05 5.39 2.24
FinTabNet 11.80 11.93 4.36 1.01
PubTables-1M 3.78 13.41 5.46 3.01
TASERTab 11.00 53.70 6.36 2.67
Table 6: Complexity of table instances across datasets. TASERTab exhibits almost five times the number of rows compared to other datasets. The maximum row count in TASERTab is 426 rows across 44 pages for a single Financial Holdings Table.
# Unique Cell Avg. # Tables Avg. Rows Avg. Columns Maximum Page Span
Dataset # Tables Topologies Per Topology Per Table Per Table
Financial Holdings Table 1933 621 3.11 53.7 6.36 44
All Other Tables 1280 331 4.32 26.9 3.87 37
Table 7: Complexity of Financial Holdings Tables
Appendix G Schema Definitions and Portfolio Model
Portfolio Base Model:
The core Instrument base model in our Pydantic model is subclassed into the following classes (see Figure 13 for the full class diagram). This is our initial schema composed of some of the most well-known financial instruments:
[…]
    Equity: a share of ownership in a corporation, representing residual claims on earnings and assets.
[…]
    Bond: a fixed‐income security issued by governments or corporations, paying periodic coupons and returning principal at maturity.
[…]
    Future: an exchange‐traded contract obligating the buyer or seller to transact an asset at a predetermined price on a specified future date.
[…]
    Forward: an over‐the‐counter agreement to buy or sell an underlying asset at a set price on a future date, customizable but counterparty‐risky.
[…]
    Swap: a bilateral contract to exchange cash flows (e.g., fixed vs. floating interest rates or different currencies), with terms set at initiation.
[…]
    Option: a derivative granting the right, but not the obligation, to buy (call) or sell (put) an underlying asset at a specified strike price before or at expiry.
[…]
    Debt: a broad class of fixed-income securities including variable return notes, medium-term notes, and government bonds, not otherwise classified as standard bonds.
[…]
    Equity Linked Note (ELN): a structured product whose returns are linked to the performance of an underlying equity or basket of equities.
[…]
    Other: a catch-all for instrument types not covered by the above classes, enabling schema extension and novelty detection.
Figure 10: Coverage vs. Utilization by Batch Size. The number of unmatched holdings covered (bars, left axis) decreases with increasing batch size, while the fraction of schema suggestions utilized (line, right axis) increases. This highlights a tradeoff: small batches are more exhaustive in coverage, but large batches are more efficient—yielding fewer “wasted” schema suggestions.
Batch Name Diversity Schema Diversity
Size Avg Min Max Avg Min Max
10 25.94 0 82 331.80 0 1387
50 22.67 0 78 313.41 0 1305
100 24.21 0 71 350.32 0 1569
250 22.60 0 55 342.97 0 1230
500 20.35 0 54 246.40 0 737
Table 8: Diversity metrics of unique schema suggestions for varying batch sizes. We report the average/minimum/maximum pairwise Levenshtein distance; “schema” metrics are over the entire generated schema, “name” is on the generated holding class name.
Appendix H Ablation Strategies
Raw Text Prompting.
For the baseline ablation, we prompt the LLM solely with the raw page text, asking whether a portfolio table is present via a simple yes/no detection prompt. Upon affirmative detection, the LLM is instructed to extract a portfolio table from the same text, returning the result as a structured object with a portfolio field, but without access to any schema or structural guidance. This strategy measures the LLM’s extraction performance in the absence of schema scaffolding or explicit reasoning.
Structured Chain-of-Thought (CoT).
To assess the impact of explicit reasoning on table detection, we prompt the LLM with the page text and require a structured Pydantic output containing both a chain-of-thought explanation (table_chain_of_thought) and a boolean indicating the presence of a portfolio table (has_portfolio_table). This ablation isolates the effect of minimal schema guidance and encourages the model to make its decision transparent through explicit intermediate reasoning. Upon positive detection, extraction is performed identically to the baseline, without additional schema context.
Full Schema Prompting.
In this ablation, we inject the complete Portfolio Pydantic schema directly into the detection prompt, alongside the page text. The LLM is instructed to reason about the presence of a portfolio table, outputting a chain-of-thought (chain_of_thought), a boolean detection (has_portfolio_table), and, if present, an extracted portfolio object conforming to the provided schema. This strategy evaluates the effect of strong schema supervision on both detection and extraction performance, requiring the model to both reason and map raw text into the structured schema within a single step.
Direct Schema Application.
For the final ablation, we bypass explicit table detection and directly apply the Portfolio schema extraction to every page. The LLM is prompted to extract a portfolio table from the provided text and return a Pydantic object with a portfolio field, irrespective of any prior detection or reasoning. Extraction is considered successful if any portion of the schema can be instantiated from the text. This approach evaluates schema-constrained extraction in the absence of explicit detection or intermediate supervision.
Appendix I Aggregation and Conflict Resolution of Schema Suggestions
After the LLM returns a batch of schema suggestions, we aggregate and cluster similar proposals as follows:
  1. 1.
    Deduplication: Suggestions with Levenshtein similarity $\geq 0.9$ (on class name and field structure) are merged.
  2. 2.
    Clustering: All proposals are clustered by semantic similarity of class names and required fields, using LLMs as the decision process.
  3. 3.
    Selection: For each cluster, the most frequent or most comprehensive schema suggestion is selected.
  4. 4.
    Validation: Each selected schema is validated by re-extracting unmatched holdings; suggestions that do not match any holding are dropped.
  5. 5.
    Manual review: If ambiguity remains, a manual review is triggered for final decision. We validated 64 resolved schemas for the second phase of extraction. Listing 15 displays the reconciled JSON schema for Forward Currency Contract, corresponding Pydantic model via pydantic.create_model, and several re-extracted holdings.
Appendix J Schema Suggestion Diversity
Table 8 summarizes the diversity among schema suggestions across batch sizes. Moderate batch sizes (100–250) achieve the highest average and maximum diversity, while the largest batch size (500) yields the lowest. This indicates that extremely large batches tend to generate more homogeneous or redundant suggestions, while moderate batches foster a broader range of candidate schemas.
Figure 11: Left: Name diversity (average, minimum, and maximum pairwise Levenshtein distance) among schema suggestions for varying batch sizes. Right: Schema diversity for the same. Moderate batch sizes (100–250) maximize diversity, while very large batches yield more homogeneous outputs.
Appendix K Example Holdings Tables
We show example holdings tables, alongside TASER’s extractions in Figures 19 - 37.
Batch Coverage (Holdings) Utilization (Schemas) Reported Collision
Size Count % Covered NAV (%) Total # # Utilized % Collisions Rate (%)
10 16,942 96.1 99.65 867 251 29.0 2,409 73.5
50 7,416 42.1 35.35 586 240 41.0 442 57.0
100 7,311 41.5 35.39 495 217 43.8 218 30.6
250 6,955 39.5 35.46 351 156 44.4 75 17.6
500 6,349 36.0 35.10 184 109 59.2 30 14.0
Table 9: Schema Utilization Efficiency. We report the proportion of generated schema suggestions that were utilized (i.e., matched at least one holding), for matching. Larger batch sizes result in a higher fraction of utilized schemas, suggesting that bulkier suggestion rounds are more efficient at targeting actionable schemas, albeit at the expense of overall diversity and coverage.
Detection Extraction End to End
Method Tokens Latency (s) Tokens Latency (s) Tokens Latency (s)
TASER (a) Raw Text Prompting 1,495 0.33 5,414 20.37 6,909 20.69
(b) Structured CoT 1,514 1.70 5,440 20.20 6,954 21.90
(c) Full Schema Prompting 5,706 1.58 5,235 20.48 10,941 22.07
(d) Direct Schema Application 5,693 21.47 5,693 21.47
Table 10: Efficiency comparison of each ablation strategy. We report the token consumption and inference latency for detection, extraction, and end-to-end processing. Raw Text Prompting minimizes detection cost (1,495 tokens, 0.33 s) and achieves a total pipeline latency of 20.69 s; Structured CoT incurs additional reasoning overhead (1.70 s) with similar extraction performance; Full Schema Prompting uses the most detection tokens (5,706) but maintains comparable end-to-end latency (22.07 s); Direct Schema Application skips the detection stage entirely, applying schema validation directly in extraction. Dashes (—) indicate stages not performed by the method.
Instrument Category Count Example
Equities 28,737 Taiwan Semiconductor Manufacturing
Debt 17,105 US Treasury 4.69% 09/05/2024
Unmatched Instruments 16,822 EUR
Forwards 8,023 Bought USD Sold KRW at 0.00072513
Options 977 Written Call Unilever 4050
Futures 720 US 5 Year Bond Future
Swaps 776 Pay fixed 3.026% receive float. (1d SOFR)
ELNs 292 BNP (Laobaixing Pharm. Chain (A)) ELN 22/07/2024
Table 11: Distribution of instrument categories in the dataset, with an example for each.
1# Ablation 1: Raw Text Prompting
2detection_prompt = (
[…]
5)
6
7
8# TableDetectionResponse Pydantic Model
9class TableDetectionResponse(BaseModel):
10 table_chain_of_thought: str = Field(...,
11 description="Chain of thoughts on if the page text contains table-like content")
12 has_portfolio_table: bool = Field(...,
13 description="True if the page has a holdings table, False otherwise")
14
15
16# Ablation 2: Structured Chain-of-Thought (CoT)
17detection_prompt = (
18 "Analyze the following text and determine if it contains a portfolio table. "
19 "Provide your chain of thought and final decision in a structured output "
[…]
22)
23
24
25# Ablation 3: Full Schema Prompting
26detection_prompt = (
27 "Using the provided Portfolio JSON schema, analyze the following text and "
28 "if it can be extracted into that schema. Provide your chain of thought. "
29 "You will output a response model object including ’chain_of_thought’, "
[…]
33)
34
35
36# Ablation 4: Direct Schema Application
37detection_prompt = (
38 "Extract a portfolio table from the following text following the Portfolio schema. "
[…]
41)
Figure 12: Detection prompts for all ablation strategies. Each section is labeled with its corresponding ablation strategy.
Listing 1: Portfolio schema with all matched instrument types.
1from enum import Enum
2from typing import Optional, List, Literal
3from pydantic import BaseModel, Field
4from datetime import datetime
5
6class BaseInstrument(BaseModel):
7 cusip: Optional[str] = Field(None, description="CUSIP identifier")
8 isin: Optional[str] = Field(None,
9 description="International Securities Identification Number")
10 ticker: Optional[str] = Field(None, description="Ticker Symbol")
11 description: Optional[str] = Field(None,
12 description="Description or name of the instrument")
13 quantity: Optional[float] = Field(None, description="Number of units held")
14 market_value: Optional[float] = Field(None, description="Market value of the holding")
15
16class Equity(BaseInstrument):
17 instrument_type: Literal["Equity"] = "Equity"
18 exchange: Optional[str] = Field(None, description="Trading exchange for the equity")
19
20class Option(BaseInstrument):
21 instrument_type: Literal["Option"] = "Option"
22 underlying: Optional[str] = Field(None, description="Identifier for the underlying asset")
23 strike_price: Optional[float] = Field(None, description="Strike price of the option")
24 expiration_date: Optional[datetime] = Field(None,
25 description="Expiration date of the option")
26 option_type: Optional[str] = Field(None, description="Call or Put option")
27
28class Swap(BaseInstrument):
29 instrument_type: Literal["Swap"] = "Swap"
30 notional_amount: Optional[float] = Field(None, description="Notional amount of the swap")
31 fixed_rate: Optional[float] = Field(None,
32 description="Fixed rate component (if applicable)")
33 floating_rate_index: Optional[str] = Field(None,
34 description="Index used for floating rate leg")
35 maturity_date: Optional[datetime] = Field(None, description="Maturity date of the swap")
36 counterparty: Optional[str] = Field(None, description="The name of the counterparty")
37
38class Forward(BaseInstrument):
39 instrument_type: Literal["Forward"] = "Forward"
40 forward_price: Optional[float] = Field(None, description="Agreed forward price")
41 settlement_date: Optional[datetime] = Field(None,
42 description="Settlement date for the forward")
43
44class Future(BaseInstrument):
45 instrument_type: Literal["Future"] = "Future"
46 contract_size: Optional[int] = Field(None, description="Size of the contract")
47 expiration_date: Optional[datetime] = Field(None,
48 description="Expiration date of the future")
49
50class Debt(BaseInstrument):
51 instrument_type: Literal["Debt"] = "Debt"
52 coupon_rate: Optional[float] = Field(None,
53 description="Annual coupon rate of the debt/bond")
54 maturity_date: Optional[datetime] = Field(None,
55 description="Maturity date of the debt/bond")
56 issuer: Optional[str] = Field(None, description="Issuer of the debt/bond")
57
58class EquityLinkedNote(BaseInstrument):
59 instrument_type: Literal["Equity Linked Note"] = "Equity Linked Note"
60 issuer: Optional[str] = Field(None, description="Issuer of the ELN")
61 product: Optional[str] = Field(None, description="Underlying product of the ELN")
62 maturity_date: Optional[datetime] = Field(None, description="Maturity date of the ELN")
Listing 2: Main Portfolio Model with Unmatched (Other) Holdings class
1class Other(BaseModel):
2 description: str = Field(...,
3 description="Text of the unknown instrument.")
4 name: str = Field(...,
5 description="Suggested classification of the description or type")
6 market_value: Optional[float] = Field(None,
7 description="Market value associated with the instrument"
8 )
9
10class Portfolio(BaseModel):
11 fund_name: Optional[str] = Field(None,
12 description="Name of the fund that the portfolio belongs to")
13 value_in_thousands: bool = Field(False,
14 description="True if the market value is based on thousands")
15 equities: Optional[List[Equity]] = Field(default_factory=list,
16 description="List of equities")
17 options: Optional[List[Option]] = Field(default_factory=list,
18 description="List of options")
19 swaps: Optional[List[Swap]] = Field(default_factory=list,
20 description="List of swaps")
21 forwards: Optional[List[Forward]] = Field(default_factory=list,
22 description="List of forwards")
23 futures: Optional[List[Future]] = Field(default_factory=list,
24 description="List of futures")
25 debt: Optional[List[Debt]] = Field(default_factory=list,
26 description="List of debt instruments")
27 elns: Optional[List[EquityLinkedNote]] = Field(default_factory=list,
28 description="List of equity linked notes")
29 other_instruments: Optional[List[Other]] = Field(default_factory=list,
30 description="The list of instruments that do not match any other type")
Figure 13: Class diagram of the initial Portfolio schema, showing the top‑level Portfolio containing a collection of Instrument objects, each subclassed into specific security types (Equity, Bond, Future, Forward, Swap, Option) to capture their unique attributes.
1# Prompt template for Recommender Agent, using batch size parameterization
2
3def recommender_agent_prompt(
4 portfolio_schema: dict,
5 unmatched_holdings: list,
6 batch_size: int,
7 start: int = 0,
8 previous_suggestions: list = None
9):
10 return """
11You are a schema refinement assistant for financial tables. Your task is:
12- Review a batch of {batch_size} unmatched financial holdings.
13- Given the current schema (JSON below), propose new classes or modifications so each holding
14 can be classified.
15- If a holding matches a previously suggested class, propose new optional fields if needed.
16- Return your schema suggestions as a list of Pydantic SchemaSuggestion model objects.
17
18Current Portfolio Schema:
19{Portfolio.model_json_schema()}
20
21Batch of unmatched holdings:
22{unmatched_holdings[start : start + batch_size]}
23
24Previously seen suggestions (optional, from prior batches):
25{previous_suggestions if previous_suggestions else None}
26
27For each unique holding, propose:
28- A new schema class, or a modification to an existing class (add or refine fields).
29- Specify all required and optional fields with Python type hints.
30- If similar to an earlier suggestion, mark only new fields as optional.
31- Provide a sample match (the original holding string).
32- Output format: a Python list of SchemaSuggestion objects, as defined below.
33"""
34
35
36class SchemaSuggestion(BaseModel):
37 name: str # Name of new or modified schema class
38 suggested_schema: str # JSON schema for the instrument.
39 example: str # Example instrument seen in unmatched holdings
Figure 14: Recommender Agent schema suggestion prompt, output model, and example LLM response. The agent sees a batched portion of unmatched holdings to recommend new alterations to the Portfolio schema. This prompt is batch-specific and may include previous_suggestions for cross-batch refinement and de-duplication.
Listing 3: Currency Forward Generated JSON Schema
1{
2 "title": "Currency Forward",
3 "type": "object",
4 "properties": {
5 "description": {
6 "type": "string",
7 "title": "Description",
8 "description": "Description or name of the currency forward"
9 },
10 "market_value": {
11 "anyOf": [
12 { "type": "number" },
13 { "type": "null" }
14 ],
15 "title": "Market Value",
16 "description": "Market value of the currency forward",
17 "default": null
18 },
19 "instrument_type": {
20 "type": "string",
21 "title": "Instrument Type",
22 "const": "Currency Forward",
23 "default": "Currency Forward"
24 },
25 "currency_pair": {
26 "anyOf": [
27 { "type": "string" },
28 { "type": "null" }
29 ],
30 "title": "Currency Pair",
31 "description": "Currency pair involved in the forward contract",
32 "default": null
33 },
34 "forward_rate": {
35 "anyOf": [
36 { "type": "number" },
37 { "type": "null" }
38 ],
39 "title": "Forward Rate",
40 "description": "Agreed forward rate",
41 "default": null
42 },
43 "settlement_date": {
44 "anyOf": [
45 { "type": "string", "format": "date-time" },
46 { "type": "null" }
47 ],
48 "title": "Settlement Date",
49 "description": "Settlement date for the currency forward",
50 "default": null
51 }
52 }
53}
Listing 4: Currency Forward Pydantic Model
1class CurrencyForward(BaseModel):
2 description: str
3 market_value: Optional[float]
4 instrument_type: str = "Currency Forward"
5 currency_pair: Optional[str]
6 forward_rate: Optional[float]
7 settlement_date: Optional[datetime]
Listing 5: Refined Extraction
1# Raw inputs
2"Bought EUR Sold USD at 0.93035372 11/06/2024"
3"Bought USD Sold GBP at 1.25473636 31/05/2024"
4"Bought GBP Sold USD at 0.79368122 16/05/2024"
5
6# Extracted as fields
7{
8 "description": "Bought EUR Sold USD at 0.93035372 11/06/2024",
9 "market_value": -282515.0,
10 "instrument_type": "Currency Forward",
11 "currency_pair": "EUR/USD",
12 "forward_rate": 0.93035372,
13 "settlement_date": "2024-06-11T00:00:00"
14},
15{
16 "description": "Bought USD Sold GBP at 1.25473636 31/05/2024",
17 "market_value": 20651.0,
18 "instrument_type": "Currency Forward",
19 "currency_pair": "USD/GBP",
20 "forward_rate": 1.25473636,
21 "settlement_date": "2024-05-31T00:00:00"
22},
23{
24 "description": "Bought GBP Sold USD at 0.79368122 16/05/2024",
25 "market_value": 1429313.0,
26 "instrument_type": "Currency Forward",
27 "currency_pair": "GBP/USD",
28 "forward_rate": 0.79368122,
29 "settlement_date": "2024-05-16T00:00:00"
30}
Figure 15: Left: Final Currency Forward JSON schema. Top right: Equivalent Pydantic model. Bottom right: Example input string and its extraction into schema fields. This demonstrates schema-driven parsing of text into structured portfolio data. A currency forward contract is a financial instrument in the foreign exchange market that locks in the price at which an entity can buy or sell a currency at a future date.
Figure 16: Holdings Table Example 1
Market Coupon Maturity
Description Quantity Value Type Rate Date Issuer Debt Type
Poland Government 0.25% 25/10/2026 329000 58000 Debt 0.25 10/25/2026 Poland Government Government Bond
Poland Government 1.25% 25/10/2030 806000 127000 Debt 1.25 10/25/2030 Poland Government Government Bond
Poland Government 1.75% 25/04/2032 822000 127000 Debt 1.75 04/25/2032 Poland Government Government Bond
Poland Government 2.25% 25/10/2024 1688000 329000 Debt 2.25 10/25/2024 Poland Government Government Bond
Poland Government 5.75% 25/04/2029 499000 103000 Debt 5.75 04/25/2029 Poland Government Government Bond
Romania Government 3.25% 24/06/2026 1035000 168000 Debt 3.25 06/24/2026 Romania Government Government Bond
Romania Government 3.65% 28/07/2025 175000 30000 Debt 3.65 07/28/2025 Romania Government Government Bond
Romania Government 3.65% 24/09/2031 610000 89000 Debt 3.65 09/24/2031 Romania Government Government Bond
Romania Government 4.75% 24/02/2025 590000 101000 Debt 4.75 02/24/2025 Romania Government Government Bond
Romania Government 6.7% 25/02/2032 415000 74000 Debt 6.7 02/25/2032 Romania Government Government Bond
Republic of South Africa 8% 31/01/2030 3362492 133000 Debt 8 01/31/2030 Republic of South Africa Government Bond
Republic of South Africa 8.5% 31/01/2037 6431353 216000 Debt 8.5 01/31/2037 Republic of South Africa Government Bond
Republic of South Africa 8.75% 31/01/2044 3011713 96000 Debt 8.75 01/31/2044 Republic of South Africa Government Bond
Republic of South Africa 8.75% 28/02/2048 3687306 116000 Debt 8.75 02/28/2048 Republic of South Africa Government Bond
Republic of South Africa 9% 31/01/2040 3143167 105000 Debt 9 01/31/2040 Republic of South Africa Government Bond
Thailand Government 2% 17/06/2042 3761000 74000 Debt 2 06/17/2042 Thailand Government Government Bond
Thailand Government 2.125% 17/12/2026 15929000 363000 Debt 2.125 12/17/2026 Thailand Government Government Bond
Thailand Government 3.3% 17/06/2038 3760000 90000 Debt 3.3 06/17/2038 Thailand Government Government Bond
Thailand Government 3.4% 17/06/2036 3668000 89000 Debt 3.4 06/17/2036 Thailand Government Government Bond
Thailand Government 3.775% 25/06/2032 4334000 108000 Debt 3.775 06/25/2032 Thailand Government Government Bond
Thailand Government 4.85% 17/06/2061 836000 23000 Debt 4.85 06/17/2061 Thailand Government Government Bond
UK Treasury 0% 08/01/2024 7000000 6994000 Debt 0 01/08/2024 UK Treasury Government Bond
UK Treasury 0% 22/01/2024 9500000 9473000 Debt 0 01/22/2024 UK Treasury Government Bond
UK Treasury 0% 19/02/2024 2900000 2880000 Debt 0 02/19/2024 UK Treasury Government Bond
UK Treasury 0% 29/04/2024 8200000 8062000 Debt 0 04/29/2024 UK Treasury Government Bond
US Treasury 4.125% 31/07/2028 13507300 10705000 Debt 4.125 07/31/2028 US Treasury Government Bond
US Treasury 4.5% 15/11/2033 19261400 15862000 Debt 4.5 11/15/2033 US Treasury Government Bond
Figure 17: Debt Extracted
Market Expiration
Description Quantity Value Type Date
CBT US 10 Year Ultra Future March 2024 181 236000 Future 03/01/2024
CBT US Ultra Bond (CBT) March 2024+ -28 0 Future 03/01/2024
EUX DAX Index Future March 2024 8 -23000 Future 03/01/2024
ICF Long Gilt Future March 2024 -21 -4000 Future 03/01/2024
NYF Mini MSCI Emerging Market Future March 2024 100 130000 Future 03/01/2024
Figure 18: Futures Extracted
Market Strike Expiration Option
Description Quantity Value Type Price Date Type
S&P 500 Index Put Option 4250 February 2024 -27 -13000 Option 4250 02/01/2024 Put
S&P 500 Index Put Option 4350 March 2024 -27 -40000 Option 4350 03/01/2024 Put
S&P 500 Index Put Option 4500 February 2024 27 36000 Option 4500 02/01/2024 Put
S&P 500 Index Put Option 4600 March 2024 27 94000 Option 4600 03/01/2024 Put
Figure 19: Options Extracted
Figure 20: Holdings Table Example 2
Market
Description Quantity Value Type Exch.
Whitehaven Coal 210649 1045544 Equity AU
China Merchants Energy Shipping (A) 835101 952512 Equity CN
Transocean 186877 801702 Equity US
China Merchants Energy Shipping 221199 252299 Equity CN
Franco-Nevada (USA) 15736 1950792 Equity CA
Wheaton Precious Metals 31475 1914624 Equity CA
Baoshan Iron & Steel (A) 1049711 1033923 Equity CN
Hindalco Industries 100895 910253 Equity IN
Zijin Mining Group (H) 66000 148118 Equity CN
Agnico Eagle Mines (US) 1244 99756 Equity CA
Container Corp of India 70391 769624 Equity IN
Techtronic Industries 32500 487513 Equity HK
Taiwan Semiconductor Manufacturing 68000 2053064 Equity TW
NAVER 15346 1976673 Equity KR
Galaxy Entertainment Group 408000 2027380 Equity HK
Yum China Holdings 37955 1729989 Equity CN
Focus Media Information Technology (A) 1342200 1342677 Equity CN
ANTA Sports Products 97200 1159888 Equity CN
Focus Media Information Technology 1094891 1095280 Equity CN
Sea ADR 9468 887814 Equity SG
Yum China Holdings 5300 248116 Equity CN
Kweichow Moutai (A) 9100 2238717 Equity CN
CP ALL (F) 739800 1501788 Equity TH
Universal Robina 437420 809466 Equity PH
Indofood CBP 165500 134750 Equity ID
Kweichow Moutai (A) 100 24601 Equity CN
Bangkok Dusit Medic Service (F) 836100 777695 Equity TH
HDFC Bank 100285 2064120 Equity IN
HDFC Bank ADR 28877 1812609 Equity IN
Hong Kong Exchanges & Clearing 29100 1199421 Equity HK
AIA Group 17000 149937 Equity HK
China Overseas Land & Investment 828000 1669841 Equity HK
Figure 21: Equities Extracted
Market Maturity
Description Quantity Value Type Issuer Product Date
CITI (Focus Media Inf. Tech (A)) ELN 12/02/2026 711200 711451 Equity Linked Note CITI Focus Media Inf. Tech (A) 12/02/2026
UBS (Focus Media Inf. Tech (A)) ELN 06/02/2025 368898 369028 Equity Linked Note UBS Focus Media Inf. Tech (A) 02/06/2025
HSBC (Focus Media Inf. Tech (A)) ELN 16/05/2025 157600 157656 Equity Linked Note HSBC Focus Media Inf. Tech (A) 05/16/2026
Figure 22: ELNs Extracted
Market Forward Settlement
Description Quantity Value Type Price Date
Bought EUR Sold USD at 0.90292802 16/10/2024 408560 2806 Forward 0.90292802 10/16/2024
Bought EUR Sold USD at 0.89751876 16/10/2024 18830 16 Forward 0.89751876 10/16/2024
Bought EUR Sold USD at 0.89544132 16/10/2024 13513 -20 Forward 0.89544132 10/16/2024
Bought EUR Sold USD at 0.89300230 16/10/2024 21836 -92 Forward 0.89300230 10/16/2024
Bought USD Sold EUR at 1.10609821 16/10/2024 12489 -101 Forward 1.10609821 10/16/2024
Figure 23: Forwards Extracted
Market Contract Expiration
Description Quantity Value Type Size Date
MSCI Malaysia Index Future 20/12/2024 1590630 58275 Future 1590630 12/20/2024
IFSC Nifty 50 Index Future 31/10/2024 -674245 4628 Future 674245 10/31/2024
MSCI Thailand Index Future 20/12/2024 117630 3600 Future 117630 12/20/2024
S&P500 Emini Index Future 20/12/2024 -1448000 -25350 Future 1448000 12/20/2024
Figure 24: Futures Extracted
Market Strike Expiration Option
Description Quantity Value Type Underlying Price Date Type Ticker
Purchased Put Nvidia 95 21/03/2025 35 19250 Option Nvidia 95 03/21/2025 Put Nvidia
Purchased Put Taiwan Semic Mfg ADR 155 20/12/2024 14 7896 Option Taiwan Semic Mfg ADR 155 12/20/2024 Put Taiwan Semic Mfg ADR
Written Call Tencent Holdings 450 30/10/2024 -16 -3171 Option Tencent Holdings 450 10/30/2024 Call Tencent Holdings
Written Call Alibaba Group Holding 110 30/10/2024 -13 -4526 Option Alibaba Group Holding 110 10/30/2024 Call Alibaba Group Holding
Written Call Techtronic Industries 115 30/10/2024 -14 -5359 Option Techtronic Industries 115 10/30/2024 Call Techtronic Industries
Written Call AIA Group 65 30/10/2024 -13 -9149 Option AIA Group 65 10/30/2024 Call AIA Group
Written Call AIA Group 62.5 30/10/2024 -13 -11841 Option AIA Group 62.5 10/30/2024 Call AIA Group
Written Call NVIDIA 125 21/03/2025 -35 -58100 Option NVIDIA 125 03/21/2025 Call NVIDIA
Figure 25: Options Extracted
Figure 26: Holdings Table Example 3
Market
Description Value Type
USD/CHF 0.1 million -1467.11 Forward
USD/DKK 0.4 million -627.32 Forward
USD/GBP 0.1 million -1408.9 Forward
USD/JPY 0.6 million -167.24 Forward
USD/NOK 1.6 million -6901 Forward
USD/SEK 0.2 million -505.44 Forward
Figure 27: Forwards Extracted
Market
Description Type Value
Cash at bank Other 1844776
Demand deposits at Depositary - EUR deposits Other 534181.49
Deposits in other EU/EEA currencies - Danish krone Other 74061.35
Deposits in other EU/EEA currencies - Norwegian krone Other 186487.92
Deposits in other EU/EEA currencies - Swedish krona Other 74121.02
Deposits in non-EU/EEA currencies - British pound Other 214117.1
Deposits in non-EU/EEA currencies - Hong Kong dollar Other 2061.23
Deposits in non-EU/EEA currencies - Japanese yen Other 2557.6
Deposits in non-EU/EEA currencies - Canadian dollar Other 81066.58
Deposits in non-EU/EEA currencies - Swiss franc Other 74577.41
Deposits in non-EU/EEA currencies - U.S. dollar Other 601544.3
Dividends/Distributions receivable Other 300071.97
Prepaid placement fee Other 879371.08
Receivables from exceeding the expense cap Other 22774.37
Other receivables Other 2326.23
Receivables from share certificate transactions Other 42598.6
Liabilities from cost items Other -566947.6
Liabilities from share certificate transactions Other -743025.51
Figure 28: Other Instruments Extracted
Figure 29: Holdings Table Example 4
Market
Description Value Type
Novo Nordisk 41930000 Equity
ASML 34798000 Equity
SAP 31184000 Equity
TotalEnergies 24622000 Equity
Siemens 23147000 Equity
UniCredit 23012000 Equity
Deutsche Boerse 19999000 Equity
Munich Re 19512000 Equity
Anheuser-Busch InBev 18885000 Equity
CRH 18877000 Equity
Sanofi 18857000 Equity
BNP Paribas 17815000 Equity
Schneider Electric 17192000 Equity
Novartis 17036000 Equity
Alcon 16833000 Equity
Safran 16448000 Equity
National Grid 16195000 Equity
Airbus 15578000 Equity
Compass 15493000 Equity
LVMH Moët Hennessy Louis Vuitton 15141000 Equity
Infineon 14905000 Equity
SGS 14327000 Equity
ASM International 13782000 Equity
Cellnex Telecom 12982000 Equity
Aena 12179000 Equity
KONE 12163000 Equity
Ryanair 11974000 Equity
Danone 11697000 Equity
Roche 11592000 Equity
Holcim 11387000 Equity
British Land 11252000 Equity
Smurfit Westrock 11087000 Equity
BAWAG 10607000 Equity
Adidas 10392000 Equity
Beiersdorf 10390000 Equity
VAT Group 10143000 Equity
Saint-Gobain 9829000 Equity
DSV 9267000 Equity
Syensqo 8991000 Equity
Anglo American 8085000 Equity
Hermès 7529000 Equity
Rheinmetall 7456000 Equity
Nestlé 7222000 Equity
Galderma 7028000 Equity
Bayer 6988000 Equity
Stellantis 5689000 Equity
Figure 30: Equities Extracted
Figure 31: Holdings Table Example 5
Description Quantity Market Value Type
Verbund 168302 11168000 Equity
Ageas 636980 24854000 Equity
Azelis Group 621052 13974000 Equity
Galapagos 139674 5397000 Equity
Umicore 526316 16565000 Equity
Novozymes B 335786 13970000 Equity
Fortum 2175101 28338000 Equity
Neste 495924 19882000 Equity
Outokumpu 4402518 19859000 Equity
Carrefour 981701 14413000 Equity
Danone 395995 17454000 Equity
Pernod Ricard 97294 16726000 Equity
Societe Generale 885047 19377000 Equity
Ubisoft Entertainment 666784 12290000 Equity
Worldline 692255 25562000 Equity
Bayer 829596 42218000 Equity
Beiersdorf 307918 29865000 Equity
Fresenius 661906 16613000 Equity
GEA Group 621022 22067000 Equity
Knorr-Bremse 321638 17138000 Equity
MTU Aero Engines 137059 27000000 Equity
Porsche Automobil Holding Preference 510872 24911000 Equity
Siemens Energy 1001433 16767000 Equity
Software 473266 10792000 Equity
Wacker Chemie 93580 11105000 Equity
Zalando 481447 18162000 Equity
Bank of Ireland Group 2251106 18185000 Equity
Intesa Sanpaolo 6737795 13635000 Equity
ASM International 38474 9953000 Equity
BE Semiconductor Industries 533800 30856000 Equity
CNH Industrial 261421 3714000 Equity
Koninklijke Philips 1316703 18517000 Equity
OCI 87808 2398000 Equity
QIAGEN 879188 36463000 Equity
Aker BP 438070 11056000 Equity
Mowi 1093259 15766000 Equity
Amadeus IT Group 266403 13491000 Equity
CaixaBank 6213097 21527000 Equity
AAK 1263617 18364000 Equity
Billerud 1817978 18498000 Equity
Elekta B 2077018 11364000 Equity
Munters Group 1521430 12313000 Equity
Mycronic 492329 7752000 Equity
Saab B 522982 16880000 Equity
SKF B 1870356 27343000 Equity
Svenska Handelsbanken A 2919741 25123000 Equity
Tele2 B 2424889 17979000 Equity
Viaplay Group B 550681 9641000 Equity
Cie Financiere Richemont 408985 49902000 Equity
Novartis 609263 45498000 Equity
Swiss Re 253533 20755000 Equity
Figure 32: Equities Extracted
Description Type Market Value
Net other assets Other 12016000
Figure 33: Other Instruments Extracted
Figure 34: Holdings Table Example 6
Market Coupon Maturity
Description Quantity Value Type Rate Date Issuer
CBT US 10 Year Ultra Future March 2024 181 236000 Future 4250 02/01/2024 Poland Government
CBT US Ultra Bond (CBT) March 2024+ -28 0 Future 4250 03/01/2024 Poland Government
EUX DAX Index Future March 2024 8 -23000 Future 4350 03/01/2024
ICF Long Gilt Future March 2024 -21 -4000 Future 4500 02/01/2024
NYF Mini MSCI Emerging Market Future March 2024 100 130000 Future 4600 03/01/2024
Figure 35: Debt Extracted
Market Expiration
Description Quantity Value Type Date
US 10 Years Note Short futures contracts -35 55508 Future 02/03/2025
US 5 Years Note Short futures contracts -17 11953 Future 31/03/2025
Figure 36: Futures Extracted
Market Type Settlement
Description Quantity Value Type Date
USD/CAD Forward Contract 1 53468 Forward 31/01/2025
USD/EUR Forward Contract 1 482874 Forward 31/01/2025
USD/GBP Forward Contract 1 43213 Forward 31/01/2025
Figure 37: Forwards Extracted